### v1

# Empirical Bernstein and betting confidence intervals for randomized quasi-Monte Carlo

Abstract
Randomized quasi-Monte Carlo (RQMC) methods estimate the mean of a random variable by sampling an integrand at $n$ equidistributed points. For scrambled digital nets, the resulting variance is typically $\tilde{O}(n^{-\theta})$ where $\theta\in[1,3]$ depends on the smoothness of the integrand and $\tilde{O}$ neglects logarithmic factors. While RQMC can be far more accurate than plain Monte Carlo (MC) it remains difficult to get confidence intervals on RQMC estimates. We investigate some empirical Bernstein confidence intervals (EBCI) and hedged betting confidence intervals (HBCI), both from Waudby-Smith and Ramdas (2024), when the random variable of interest is subject to known bounds. When there are $N$ integrand evaluations partitioned into $R$ independent replicates of $n=N/R$ RQMC points, and the RQMC variance is $\Theta(n^{-\theta})$, then an oracle minimizing the width of a Bennett confidence interval would choose $n=\Theta(N^{1/(\theta+1)})$. The resulting intervals have a width that is $\Theta(N^{-\theta/(\theta+1)})$. Our empirical investigations had optimal values of $n$ grow slowly with $N$, HBCI intervals that were usually narrower than the EBCI ones, and optimal values of $n$ for HBCI that were equal to or smaller than the ones for the oracle.
1 Introduction
We study the combination of randomized quasi-Monte Carlo (RQMC) integration to estimate an expectation with some non-asymptotic (finite sample valid) methods to get a confidence interval for such an expectation. RQMC has significant accuracy benefits over plain Monte Carlo (that we outline below), but the usual confidence intervals based on RQMC estimates have justifications that are asymptotic in their sample size LEcEtal24a . Empirical Bernstein and some related betting methods (described below) give finite sample assurances for bounded integrands. Specifically, if it is known that the integrand values must lie between 00 and $1$, then those methods can construct a confidence interval that has at least $1-\alpha$ probability to contain the integral’s value. This confidence level holds for any integrand satisfying the given bounds.
The context for our work is as follows. Many scientific problems require the numerical evaluation of multidimensional integrals. When the dimension is high enough then classical tensor product integration rules like those in DavRab84 become too expensive to use. It is then common to use plain Monte Carlo (MC) methods instead. MC methods based on random sampling typically have root mean squared errors (RMSEs) of $O(n^{-1/2})$ given $n$ function evaluations. This rate is slow but it is the same in any dimension. MC also allows one to use statistical methods to quantify the uncertainty in the estimated integral. Quasi-Monte Carlo (QMC) methods (e.g., DicPil10a and Nie92 ) are deterministic sampling strategies that under a bounded variation assumption produce integral estimates with error $\tilde{O}(n^{-1})$ where $\tilde{O}(\cdot)$ means we neglect powers of $\log(n)$. Plain QMC methods are deterministic and so they lose the uncertainty quantification advantage of MC. RQMC methods, surveyed in LEcLem02a , allow one to make independent replicates of a statistically unbiased QMC method to support variance estimates and asymptotic confidence intervals. RQMC methods can also improve on the convergence rate of QMC methods obtaining an RMSE of $\tilde{O}(n^{-3/2})$ under a smoothness assumption Owe97 ; Owe08a on the integrand.
In addition to asymptotic confidence interval methods for RQMC, some other papers have provided finite sample uncertainty quantifications based on additional assumptions. For purely independent and identically distributed (IID) sampling, knowing a bound on the kurtosis facilitates a finite sample confidence interval for the mean of a random variable in terms of the sample mean and the sample standard deviation HicEtal14a . For deterministic QMC sampling of $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ with lattices or digital nets, assuming that the coefficients in a suitable orthogonal decomposition of $f:[0,1]^{d}\to\mathbb{R}$ decay reasonably, one may derive a deterministic error bound on $\mu=\mathbb{E}[f(\boldsymbol{x})]$ HicJim16a ; HicEtal17a ; JimHic16a ; JagSor23a . A different approach is to assume that $f$ is a realization of a Gaussian process and construct a credible interval for $\mathbb{E}[f(\boldsymbol{x})]$. If the covariance kernel and the sampling nodes are well-matched, the computational effort required is nearly $O(n)$ and not the $O(n^{3})$ effort typically required for such credible intervals RatHic19a ; JagHic22a . The GAIL Gail_ug and QMCPy QMCPy2020a libraries implement these finite sample uncertainty quantifications under the assumptions outlined above.
Here, we study RQMC confidence intervals assuming that the integrand is bounded between zero and one. The advantage of this assumption is that there are more settings where we are certain that it holds.
The non-asymptotic confidence intervals that we emphasize are the recently developed hedged betting confidence intervals (HBCI) and some related predictable plug-in empirical Bernstein confidence intervals (EBCI), both from WauRam24a . They are derived from some infinite confidence sequences and we believe that they provide the narrowest confidence intervals among currently available methods.
Section 2 gives our notation and some definitions and background on quasi-Monte Carlo, randomized quasi-Monte Carlo, and the EBCI and HBCI of WauRam24a . In Section 3 we show how splitting $N$ observations into $R=N/n$ independent replicates of $n$ RQMC point sets can give narrower EBCI than MC does for an oracle that knows the RQMC variance at each $n$. The extent of the narrowing depends on the RQMC convergence rate which varies from problem to problem. Perhaps surprisingly, the more effective RQMC is, the smaller the optimal value of $n$ becomes. For an RQMC variance proportional to $n^{-\theta}$, taking $n$ proportional to $N^{1/(\theta+1)}$ gives the narrowest intervals. The resulting interval widths are $\Theta(N^{-\theta/(\theta+1)})$ which can be much wider than the assumed RQMC standard deviation of $\Theta(N^{-\theta/2})$. Section 4 makes some computational investigations. As predicted by the theory the best values of $n$ to use grew slowly with $N$. The empirically best values of $n$ for HBCI were no larger than the oracle values, and sometimes smaller. Section 5 has some final comments and discussion.
2 Background and notation
We begin with some notation. We use $1\{E\}$ to denote a quantity that equals $1$ when expression $E$ holds and is zero otherwise. The vectors $\boldsymbol{0}$ and $\boldsymbol{1}$ have all components equal to 00 and $1$ respectively with a dimension given by context. We write $a_{n}=O(b_{n})$ as $n\to\infty$ if there exists $C>0$ and $n_{0}<\infty$ with $|a_{n}|\leqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Omega(b_{n})$ if there exists $C>0$ and $n_{0}<\infty$ with $a_{n}\geqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$.
2.1 Intervals and martingales
For $0\leqslant\alpha\leqslant 1$, a $1-\alpha$ confidence interval for a parameter $\mu$ is a pair of random quantities $A$ and $B$ with
$\Pr(A\leqslant\mu\leqslant B)\geqslant 1-\alpha.$ (1)
A confidence interval is strict when the right hand side above is an equality. An asymptotic confidence interval is a sequence $(A_{n},B_{n})$ of random quantities for which
$\lim_{n\to\infty}\Pr(A_{n}\leqslant\mu\leqslant B_{n})\geqslant 1-\alpha.$
It is common for asymptotic confidence intervals to be strict, meaning that this limit equals $1-\alpha$. The best known example is Student’s $t$ confidence interval for the mean of a distribution with finite variance, based on a random sample from that distribution and justified by the central limit theorem.
The random sequence $(A_{n},B_{n})$ is a $1-\alpha$ confidence sequence for $\mu$ if
$\Pr(A_{n}\leqslant\mu\leqslant B_{n},\ \forall n\geqslant 1)\geqslant 1-\alpha.$ (2)
Confidence sequences allow us to select a stopping time $\nu$ based on $\{(A_{n},B_{n})\mid n\leqslant\nu\}$ and have at least $1-\alpha$ confidence that $A_{\nu}\leqslant\mu\leqslant B_{\nu}$. We can take the limits to be $\max_{n\leqslant\nu}A_{n}$ and $\min_{n\leqslant\nu}B_{n}$.
Martingale theory is one of the main tools for obtaining confidence sequences. If random variables $S_{i}$ for $i\geqslant 1$ satisfy $\mathbb{E}(S_{n}\!\mid\!S_{1},\dots,S_{n-1})=S_{n-1}$ then $(S_{i})_{i\geqslant 1}$ is a martingale. If $\mathbb{E}(S_{n}\!\mid\!S_{1},\dots,S_{n-1})\leqslant S_{n-1}$ then $(S_{i})_{i\geqslant 1}$ is a supermartingale. Ville’s inequality is that for any $\eta>0$ and any nonnegative supermartingale $(S_{i})_{i\geqslant 1}$
$\Pr\biggl{(}\sup_{n\geqslant 1}S_{n}\geqslant\eta\biggr{)}\leqslant\frac{\mathbb{E}(S_{1})}{\eta}.$
The confidence sequences from WauRam24a are derived by applying Ville’s inequality vill:1939 to nonnegative supermartingales.
The problem we consider is to approximate the finite-dimensional integral
$\mu=\int_{[0,1]^{d}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}=\mathbb{E}(f(\boldsymbol{x})),\quad\text{for $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$}.$
In MC sampling we take $\boldsymbol{x}_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathbb{U}[0,1]^{d}$ and estimate $\mu$ by
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}f(\boldsymbol{x}_{i}).$
Let $\sigma^{2}=\operatorname{var}(f(\boldsymbol{x}))$ for $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$. We assume that $0<\sigma^{2}<\infty$. In that case the RMSE of $\hat{\mu}$ is $\sigma/\sqrt{n}$. For $n\geqslant 2$,
$s^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(f(\boldsymbol{x}_{i})-\hat{\mu})^{2}$
is an unbiased estimate of $\sigma^{2}$. Furthermore
$\lim_{n\to\infty}\Pr\Bigl{(}\sqrt{n}\frac{\hat{\mu}-\mu}{s}\leqslant t^{1-\alpha}_{(n-1)}\Bigr{)}=1-\alpha$
where $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on $n-1$ degrees of freedom. Then $\hat{\mu}\pm st^{1-\alpha/2}/\sqrt{n}$ is an asymptotic $1-\alpha$ confidence interval for $\mu$.
2.2 QMC and RQMC
QMC sampling replaces random points $\boldsymbol{x}_{i}$ by deterministic points that more evenly sample the unit cube. There are many ways to quantify that property. The most elementary one is the star discrepancy:
$D_{n}^{*}=D_{n}^{*}(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})=\sup_{\boldsymbol{a}\in[0,1]^{d}}\,\biggl{|}\frac{1}{n}\sum_{i=1}^{n}1\{\boldsymbol{x}_{i}\in[\boldsymbol{0},\boldsymbol{a})\}-\prod_{j=1}^{d}a_{j}\biggr{|}.$
Small values of $D_{n}^{*}$ show that all anchored boxes $[\boldsymbol{0},\boldsymbol{a})$ have very nearly the desired proportion of the $n$ points, which is the volume of $[\boldsymbol{0},\boldsymbol{a})$. It is possible to choose points $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}$ so that $D_{n}^{*}=\tilde{O}(n^{-1})$ while plain MC points have $D_{n}^{*}=\tilde{O}(n^{-1/2})$ Nie92 .
The improvement in star discrepancy yields an improvement in integration error via the Koksma-Hlawka inequality Hic04a :
$|\hat{\mu}-\mu|\leqslant D_{n}^{*}(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})\times V_{\mathrm{HK}}(f).$ (3)
As described above $D_{n}^{*}$ is a measure of how non-uniform the points $\boldsymbol{x}_{i}$ are. The new factor $V_{\mathrm{HK}}(f)$ is the total variation of $f$ in the sense of Hardy and Krause. See Owe05a . It follows from (3) that we can get $|\hat{\mu}-\mu|=\tilde{O}(n^{-1})$ using QMC. We will refer to this as the BVHK rate.
If we knew $D_{n}^{*}$ and $V_{\mathrm{HK}}(f)$, then (3) would provide a perfect quantification of our uncertainty about $\mu$. It would be non-asymptotic and even non-probabilistic. Unfortunately, $D_{n}^{*}$ is generally very expensive to obtain and $V_{\mathrm{HK}}(f)$ is ordinarily far harder to compute than $\mu$. As a result, equation (3) shows us that MC can be outperformed but it does not provide a generally usable error estimate.
RQMC methods generate random points $\boldsymbol{x}_{i}$ that simultaneously satisfy two properties:
  1. 1)
    $\boldsymbol{x}_{i}\sim\mathbb{U}[0,1]^{d}$, for all $i=1,\dots,n$, and
  2. 2)
    $D_{n}^{*}(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})=\tilde{O}(n^{-1})$ almost surely.
From the second property, $|\hat{\mu}-\mu|=\tilde{O}(n^{-1})$. As a result the RMSE of RQMC is $\tilde{O}(n^{-1})$ in the BVHK case.
To get an approximate confidence interval from RQMC we can form $R$ statistically independent RQMC estimates $\hat{\mu}_{1},\dots,\hat{\mu}_{R}$. Using the first property above these IID estimates are unbiased and we can get an asymptotic confidence interval for $\mu$ of the form
$\hat{\mu}\pm St_{(R-1)}^{1-\alpha/2}/\sqrt{R}$
for
$\hat{\mu}=\frac{1}{R}\sum_{i=1}^{R}\hat{\mu}_{i}\quad\text{and}\quad S^{2}=\frac{1}{R-1}\sum_{i=1}^{R}(\hat{\mu}_{i}-\hat{\mu})^{2}.$
Here $S^{2}/R$ is an unbiased estimate of the MSE of the pooled estimate $\hat{\mu}$. The confidence level $1-\alpha$ is obtained in the limit as $R\to\infty$ naka:tuff:2024 for fixed $n$.
In addition to getting an unbiased estimate of the RQMC variance, randomization can provide some other benefits. First, some RQMC methods can attain better accuracy than plain QMC. Under sufficient smoothness, the scrambled digital sequences in Owe95 attain RMSEs of $\tilde{O}(n^{-3/2})$ Owe97 ; Owe08a . The same holds for the scramblings of Mat98 . We will refer to this as the ‘smooth case’. Higher order digital nets Dic11a can attain even better convergence rates, though the rates are not necessarily evident empirically when $d$ is moderately large, which nuyens2010higher attribute to numerical precision.
Those RQMC methods have further useful properties. First, $\operatorname{var}(\hat{\mu})=o(1/n)$ as $n\to\infty$ for any integrand with $\sigma^{2}<\infty$, and so their efficiency with respect to MC becomes unbounded as $n\to\infty$. Furthermore, the condition of finite variation in the sense of Hardy and Krause can be quite strict. For example, when $d\geqslant 2$, step discontinuities in the integrand $f$ ordinarily make $V_{\mathrm{HK}}(f)$ infinite unless those discontinuities are axis parallel Owe05a . This makes QMC unattractive for estimating integrals of binary valued functions, while RQMC will still attain an RMSE that is $o(n^{-1/2})$.
While there are numerous RQMC methods in use, we will restrict ourselves here to scramblings of digital nets and sequences, such as the ones of Sobol’ Sob67 . Our variance formulas assume that the scramblings are either those of Owe95 or Mat98 . In our computations we used the linear matrix scramble of Mat98 with a digital shift.
While RQMC methods yield very good accuracy, most of the known confidence interval methods for them supply only asymptotic confidence as $R\to\infty$ naka:tuff:2024 . In some very limited settings there are central limit theorems for $\hat{\mu}$ as $n\to\infty$ Loh01 . Having the confidence statements be asymptotic in $R$ is unfortunate since the RMSE of $\hat{\mu}$ is proportional to $R^{-1/2}$ while the RMSE vanishes at a faster rate in $n$.
2.3 Hoeffding and Empirical Bernstein intervals
We would like a confidence interval that is non-asymptotic, meaning that the coverage guarantee (1) holds for limits $A$ and $B$ computed from a sample of size $n$. This can often be obtained under parametric statistical models in which the distribution of $f(\boldsymbol{x})$ belongs to a known finite dimensional family such as the Gaussian distributions. It is not desirable to make such assumptions and so we prefer a nonparametric method with non-asymptotic confidence.
There is a theorem of Bahadur and Savage BahSav56 that reveals sharp constraints on our ability to construct nonparametric and non-asymptotic confidence intervals. Here is the description from err4qmc :
They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)$ be $\mathbb{E}(Y)$ when $Y\sim F\in\mathbb{R}$, their conditions are:
  1.  (i)
    For all $F\in\mathcal{F}$, $\mu(F)$ exists and is finite.
  2.  (ii)
    For all $m\in\mathbb{R}$ there is $F\in\mathcal{F}$ with $\mu(F)=m$.
  3.  (iii)
    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $10<\pi<$ 1, then $\pi F+(1-\pi)G\in\mathcal{F}$.
Then their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in\mathbb{R}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
We can escape the restriction from Bahadur and Savage by stipulating that every distribution in $\mathcal{F}$ has support in a known interval $[a,b]$ of finite length. This violates their clause (ii). After a linear transformation we take that interval to be $[0,1]$.
We will use $Y_{i}=\hat{\mu}_{i}$ for $i=1,\dots,R$ to represent the values of the $R$ replicated RQMC points that go into our confidence interval calculation. We set $\bar{Y}_{R}=(1/R)\sum_{i=1}^{R}Y_{i}$. We assume that $Y_{i}$ are IID with the same distribution as $Y$ and that $0\leqslant Y\leqslant 1$ holds with probability one. Then
$\bar{Y}_{R}\pm\sqrt{\frac{\log(2/\alpha)}{2R}}$ (4)
for $10<\alpha<$ 1 is a $1-\alpha$ confidence interval for $\mu=\mathbb{E}(Y_{i})$. Equation (4) gives the confidence interval of Hoeffding hoef:1963 . The interval is not always nested within $[0,1]$, but we can simply take its intersection with $[0,1]$. That will preserve the confidence level.
While Hoeffding’s interval is a valid confidence interval for the mean of any distribution supported on $[0,1]$, it can be quite conservative for some of them, giving confidence over $1-\alpha$. This raises the possibility of getting coverage $1-\alpha$ from narrower intervals. One approach is to make use of $\sigma^{2}=\operatorname{var}(Y)$ if it is known. Hoeffding’s confidence interval is a consequence of Hoeffding’s exponential probability bound in hoef:1963 . Other exponential probability bounds take advantage of a known value $\sigma^{2}=\operatorname{var}(Y)$ or even a sample estimate of that variance. Those bounds can be used to get improved confidence intervals.
Theorem 3 of Maurer and Pontil maurer2009empirical is that
$\mathbb{E}(Y)-\frac{1}{R}\sum_{i=1}^{R}Y_{i}\leqslant\sqrt{\frac{2\sigma^{2}\log(1/\delta)}{R}}+\frac{\log(1/\delta)}{3R}.$
holds with probability at least $1-\delta$. They call this Bennett’s inequality, because it derives from an inequality in benn:1962 . They do not supply a proof, but it can be proved from Bernstein’s inequality given as Theorem 3 of bouc:lugo:bous:2003 . Taking $\delta=\alpha/2$ we get
$\bar{Y}_{R}\pm\sqrt{\frac{2\sigma^{2}\log(2/\alpha)}{R}}+\frac{\log(2/\alpha)}{3R}$ (5)
as a $1-\alpha$ confidence interval for $\mu$. We will make use of the Bennett confidence intervals (5) for some theoretical investigations.
In most practical settings where $\mu$ is unknown, $\sigma^{2}$ will also be unknown. Replacing $\sigma^{2}$ by $s^{2}$ in (5) would not always give a valid confidence interval. Theorem 4 of maurer2009empirical shows that for $R\geqslant 2$
$\bar{Y}_{R}\pm\sqrt{\frac{2s^{2}\log(4/\alpha)}{R}}+\frac{7\log(4/\alpha)}{3(R-1)}.$ (6)
is a $1-\alpha$ confidence interval for $\mu$. This is called an empirical Bernstein confidence interval named after an inequality of Bernstein that is similar to Bennett’s. We see that both terms on the right hand side of (6) are larger than the ones in the Bennett interval (5). Their interval has improved asymptotic constants compared to the original empirical Bernstein inequality from audi:muno:szep:2007 . The EBCI and HBCI from WauRam24a provide a further asymptotic improvement.
2.4 Betting methods
Our interest in this problem was sparked in part by some recent work by WauRam24a on confidence intervals generated by betting arguments. They consider a more general setting of confidence sequences, as in equation (2), that can then be specialized to confidence intervals as needed. For random $Y_{i}\in[0,1]$ they assume that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=\mu$ for all $i\geqslant 1$ without assuming that the $Y_{i}$ are identically distributed or even that they are independent. For $i=1$, the conditional expectation above is simply $\mathbb{E}(Y_{1})=\mu$.
Suppose that the conditional mean is really $\mu$ and there is a null hypothesis $H_{0}(m)$ that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=m$. Then if $m\neq\mu$ somebody starting with a stake of \$1.00 can make a series of bets against $H_{0}(m)$ as the $Y_{i}$ are revealed in order and have an expected fortune that grows without bound. Just prior to time $i$, the bettor picks a value $\lambda_{i}\in(-1/(1-m),1/m)$. The bettor’s capital at time $t\geqslant 1$ is
$\mathcal{K}_{t}(m)=\prod_{i=1}^{t}\bigl{(}1+\lambda_{i}(m)(Y_{i}-m)\bigr{)}.$
Taking $\lambda_{i}(m)>0$ amounts to betting that $Y_{i}>m$. Then their capital is multiplied by $1+\lambda_{i}(Y_{i}-m)$. If they are right then their capital grows, but if $Y_{i}<m$, then it shrinks. To take $\lambda_{i}(m)<0$ is to bet that $Y_{i}<m$.
Section 4 of WauRam24a notes that $\mathcal{K}_{t}(\mu)$ is a nonnegative martingale, so that $\mathbb{E}(\mathcal{K}_{t}(\mu))=1$ for all $t\geqslant 1$. There is no way for the bettor to pick bet sizes $\lambda_{i}(\mu)$ to make an expected profit. Then from Ville’s Theorem
$\Pr(\mathcal{K}_{t}(\mu)\leqslant 1/\alpha,\ \forall t\geqslant 1)\geqslant 1-\alpha.$
If we watch $\mathcal{K}_{t}(\mu)$ indefinitely, then there is at most probability $\alpha$ that it will ever go above $1/\alpha$. As a result we can get a confidence sequence by retaining at time $t$, all the values $m$ for which $\max_{1\leqslant\tau\leqslant t}\mathcal{K}_{\tau}(m)\leqslant 1/\alpha$. This idiom uses a hypothetically infinite ensemble of bettors to do this, but WauRam24a give implementable algorithms for it.
The other side of the coin is that there are good betting strategies when $m\neq\mu$. Under those, the bettor’s expected fortune will grow steadily when $m\neq\mu$. That is what one needs to make a confidence sequence not just valid but also useful in having a width that converges to zero with increased sampling.
One version of the betting strategy above produces confidence sequences that are analogous to Hoeffding intervals and another produces empirical Bernstein intervals. We present the predictable plug-in EBCI followed by a hedged betting interval that combines two betting strategies. We do not include the predictable plug-in Hoeffding intervals. Those do not use a sample variance, and they are wider than the others for larger $t$.
The confidence intervals we study make use of running sample moments
$\hat{\mu}_{t}=\frac{1/2+\sum_{i=1}^{t}Y_{i}}{t+1}\quad\text{and}\quad\hat{\sigma}^{2}_{t}=\frac{1/4+\sum_{i=1}^{t}(Y_{i}-\hat{\mu}_{i})^{2}}{t+1}$
with the means and standard deviations both biased slightly towards $1/2$. In their ‘predictable plug-in empirical Bernstein’ intervals, the betting amounts are
$\lambda_{t}^{\text{PrPl-EB}}=\sqrt{\frac{2\log(2/\alpha)}{\hat{\sigma}^{2}_{t-1}t\log(1+t)}}\,\wedge c$
for some $c\in(0,1)$ with $c=1/2$ or $3/4$ given as reasonable defaults. Note that this method only bets that $Y_{i}>m$. Theorem 2 of WauRam24a gives a $1-\alpha$ confidence sequence
$C_{t}^{\text{PrPl-EB}}=\frac{\sum_{i=1}^{t}\lambda_{i}Y_{i}}{\sum_{i=1}^{t}\lambda_{i}}\pm\frac{\log(2/\alpha)+\sum_{i=1}^{t}\nu_{i}\psi_{\mathrm{E}}(\lambda_{i})}{\sum_{i=1}^{t}\lambda_{i}},$ (7)
where
$\nu_{i}=4(Y_{i}-\hat{\mu}_{i-1})^{2}\quad\text{and}\quad\psi_{\mathrm{E}}(\lambda)=(-\log(1-\lambda)-\lambda)/4.$
The running intersection $\cap_{1\leqslant\tau\leqslant t}C^{\text{PrPl-EB}}_{\tau}$ is also a $1-\alpha$ confidence sequence for $\mu$ and of course we can intersect any of these intervals with $[0,1]$.
These confidence sequences can be specialized to confidence intervals when we have a fixed target sample size $R$ in mind. For that case WauRam24a recommends
$\lambda_{i}^{\text{PrPl-EB}(R)}=\sqrt{\frac{2\log(2/\alpha)}{R\hat{\sigma}^{2}_{i-1}}}\,\wedge c,\quad i=1,\dots,R.$ (8)
The above formula for $\lambda_{i}$ uses the ordering of the data values and puts unequal weight on the $R$ different $Y_{i}$. There is a permutation strategy in WauRam24a to treat the $Y_{i}$ values more symmetrically, but they report that it makes little difference.
The EBCI we study take the form
$\bigcap_{1\leqslant\tau\leqslant t}C_{t}^{\text{PrPl-EB}(R)}$
where $C_{t}^{\text{PrPl-EB}(R)}$ is the interval in (7) using $\lambda_{i}$ from (8). Equation (17) of WauRam24a gives the scaled half width of these EBCI as
$\sqrt{R}\biggl{(}\frac{\log(2/\alpha)+\sum_{i=1}^{R}\nu_{i}\psi_{\mathrm{E}}(\lambda_{i})}{\sum_{i=1}^{R}\lambda_{i}}\biggr{)}\stackrel{{\scriptstyle\mathrm{a.s.}}}{{\to}}\sigma\sqrt{2\log(2/\alpha)}$
as $R\to\infty$. The corresponding limit for the empirical Bernstein CIs of maurer2009empirical is $\sigma\sqrt{2\log(4/\alpha)}$.
The hedged betting process of WacMenSch08 works as follows. For positive predictable sequences $\lambda_{i}^{+}(m)$ and $\lambda_{i}^{-}(m)$ defined below, they use capital processes
$K_{t}^{+}(m)=\prod_{i=1}^{t}(1+\lambda_{i}^{+}(m)(Y_{i}-m)\quad\text{and}\quad K_{t}^{-}(m)=\prod_{i=1}^{t}(1-\lambda_{i}^{+}(m)(Y_{i}-m)$
that bet, respectively, on $Y_{i}>m$ and $Y_{i}<m$. The hedged process is
$K_{t}^{\pm}(m)=\max\bigl{(}\theta K_{t}^{+}(m),(1-\theta)K_{t}^{-}(m)\bigr{)}$
for $0\leqslant\theta\leqslant 1$. For any $\theta$, $K_{t}^{\pm}(\mu)$ is a nonnegative martingale hence subject to Ville’s theorem. The supplement to WauRam24a suggests $\theta=1/2$ as a default. If $m\neq\mu$ then $K_{t}^{\pm}(m)$ will tend to grow with $t$. Of course the hedging $\theta$ has nothing to do with the RQMC variance rate $\theta$ and the quantities are distinct enough that no confusion should arise.
The hedged betting process produces confidence sequences and, as they did for the empirical Bernstein processes, they recommend a way to produce a confidence interval based on $R$ sample values. For that they take
$\lambda_{i}^{+}(m)=|\tilde{\lambda}_{i}^{\pm}|\wedge\frac{c}{m}\quad\text{and}\quad\lambda_{i}^{-}(m)=|\tilde{\lambda}_{i}^{\pm}|\wedge\frac{c}{1-m}$
for $i=1,\dots,R$ where
$\tilde{\lambda}_{i}^{\pm}=\sqrt{\frac{2\log(2/\alpha)}{R\hat{\sigma}^{2}_{i-1}}}$
for $c\in[0,1)$, with suggested defaults of $c=1/2$ or $c=3/4$. The confidence interval is then
$\bigl{\{}m\in[0,1]\mid\max_{1\leqslant i\leqslant R}K_{i}^{\pm}(m)\leqslant 1/\alpha\bigr{\}}.$
3 Asymptotic confidence interval comparisons
Our confidence intervals will use $R$ IID replicates that each use $n$ RQMC points in $[0,1]^{d}$. That takes $N=nR$ evaluations of $f$. Here we study how to get the narrowest confidence intervals with that budget of $N$ function evaluations.
We take
$Y_{i}=\frac{1}{n}\sum_{k=1}^{n}f(\boldsymbol{x}_{i,k})$
where for each $i$, $\boldsymbol{x}_{i,1},\dots,\boldsymbol{x}_{i,n}$ is an RQMC point set. The $Y_{i}$ are IID when we use independent scrambles for all $R$ point sets. Let $\sigma^{2}_{n}=\operatorname{var}(Y_{i})$. We know that $\sigma^{2}_{n}=o(n^{-1}).$ For the BVHK case, $\sigma^{2}_{n}=\tilde{O}(n^{-2})$. For the smooth case, $\sigma^{2}_{n}=\tilde{O}(n^{-3})$.
We know from WauRam24a that the widths of their predictable plug-in EBCI approaches that of some Bennett intervals described below, as $R\to\infty$ for fixed $n$. The same is true of their HBCI. Here we study the value of $n$ that optimizes the Bennett intervals.
As mentioned above, the tilde in $\tilde{O}$ hides powers of $\log(n)$. Those powers are present in error bounds for adversarially chosen integrands but do not seem to come up in real problems schl:2002 ; wherearethelogs . We will consider working models of the form
$\sigma^{2}_{n}=\sigma^{2}_{0}n^{-\theta},$ (9)
even though the actual variance is not necessarily any power of $n$. This approach is less cumbersome than bounding the variance between $\Omega(n^{-\theta})$ and $O(n^{-\theta+\epsilon})$, and we believe it provides clearer insights.
We know that taking $\theta=1$ underestimates the quality of RQMC because the variance is $o(1/n)$ for the RQMC methods we use. The smooth case with $\theta=3$ is optimistic while the BVHK case with $\theta=2$ is intermediate.
If $R\to\infty$ for fixed $n$, then the predictable plug-in EBCI have half widths $H^{\text{PrPl-EB}}$ that satisfy
$\sqrt{R}H^{\text{PrPl-EB}(R)}\to\sigma_{n}\sqrt{2\log(2/\alpha)}.$ (10)
The HBCI have half widths $H^{\text{PrPl}}$ that satisfy the same rate. As a result, both methods give interval widths that are $O(n^{-\theta/2}R^{-1/2})$ as $R\to\infty$. For a fixed product $N=nR$, this width is narrowest at $R=1$ and $n=N$, but of course that is infeasible and also that argument ignores the fact that (10) is based on a limiting argument as $R\to\infty$.
To study the tradeoff between $n$ and $R$, we use Bennett’s inequality which gives a half width of
$H^{\mathrm{Ben}}(n,R)=\sqrt{\frac{2\sigma_{n}^{2}\log(2/\alpha)}{R}}+\frac{\log(2/\alpha)}{3R}$ (11)
when using $R$ IID observations that each have variance $\sigma^{2}_{n}$. An oracle that knew $\sigma^{2}_{n}$ could use this formula to select the best $n$ and $R$ for a confidence interval subject to a constraint $nR=N$. The predictable plug-in EBCI and the HBCI from WauRam24a do not assume a known variance for the $Y_{i}$. However those intervals attain the same asymptotic limiting width as $R\to\infty$ that an oracle would get from Bennett’s formula.
Next, we investigate the oracle’s choice of $n$ and $R$ to minimize the half width subject to $nR=N$. While $n$ and $R$ both have to be positive integers with product $N$, we will first relax the problem to a continuous variable $n$ with $R=N/n$. After that, we discuss integer solutions.
Theorem 1.
Let $\sigma^{2}_{n}$ follow (9) and choose $N>0$. If $\theta=1$, then the minimizer of (11) over $n\in[1,\infty)$ is $n_{*}=1$. If $\theta>1$, then the minimizer of (11) over $n\in(0,\infty)$ is
$n_{*}$ $=\left(\frac{9(\theta-1)^{2}\,\sigma_{0}^{2}N}{2\log(2/\alpha)}\right)^{1/(\theta+1)}.$ (12)
Proof.
Under the model (9), with $R=N/n$ the oracle’s half width is
$H(n)=H^{\mathrm{Ben}}(n,N/n)=\sigma_{0}n^{(1-\theta)/2}N^{-1/2}\sqrt{2\log(2/\alpha)}+\frac{\log(2/\alpha)}{3}\frac{n}{N}.$ (13)
If $\theta=1$, then $H(n)$ is minimized over $[1,\infty)$ at $n=1$.
For $\theta>1$, the function $H(n)$ is a convex function of $n$. It has a unique minimum over $(0,\infty)$ at some $n_{*}$ where $H^{\prime}(n_{*})=0$. Now
$H^{\prime}(n)=\frac{1-\theta}{2}\frac{\sigma_{0}}{\sqrt{N}}n^{-(\theta+1)/2}\sqrt{2\log(2/\alpha)}+\frac{\log(2/\alpha)}{3N}.$
This vanishes at
$n^{-(\theta+1)/2}=\frac{2\log(2/\alpha)/(3N)}{(\theta-1)\sigma_{0}\sqrt{2\log(2/\alpha)/N}}=\frac{\sqrt{2\log(2/\alpha)}}{3(\theta-1)\sigma_{0}\sqrt{N}}$
from which (12) follows. ∎
If $n_{*}<1$, then $n=1$ is best. This includes the degenerate case with $\sigma_{0}=0$. If $n_{*}>N$, then $n=N$ is best. If $1<n_{*}<N$ and $n_{*}$ is not an integer then the best integer $n$ is either $\lfloor n_{*}\rfloor$ or $\lceil n_{*}\rceil$ by the convexity noted in the proof.
For scrambled Sobol’ points it is generally best to take $n$ to be a power of 2 Owe22a , which is not reflected in Theorem 1. We incorporate that restriction into some numerical examples in Section 4.
As noted above, when $\theta=1$, the best $n$ is $n=1$. In that case RQMC is the same as MC, using one uniformly distributed point in $[0,1]$. This result holds for any sampler that gets the ordinary MC rate, even if it has a favorable constant. In particular, any MC variance reduction techniques that do not change the convergence rate will not lead to using $n>1$. They will improve the asymptotic confidence interval width by reducing $\sigma_{0}$.
When $\theta>1$, we see that the shortest intervals come from taking $n=\Theta(N^{1/(\theta+1)})$. The BVHK rate ($\theta=2$) is then $n=\Theta(N^{1/3})$ while the rate for smooth integrands ($\theta=3$) is $n=\Theta(N^{1/4})$. We see that when the RQMC convergence rate becomes better, the optimal RQMC sample size to use grows at a slower rate in $N$. This may seem like a paradox, but the explanation can be seen in equation (11). If $\sigma_{n}$ drops very quickly then the $O(1/\sqrt{R})$ term it is in does not dominate the other $O(1/R)$ term as much and a larger $R$ is better.
Because the oracle takes $n$ to be a smaller power of $N$ for larger $\theta$, we might wonder why it always picks $n=1$ for the smallest value, $\theta=1$. This stems from the factor $\theta-1$ in the constant of proportionality in (12). For fixed $\sigma_{0}$, $N$ and $\alpha$ we could maximize (12) over $\theta$ to see which rate gives the largest $n$.
What we see from Theorem 1 is that the optimal $n$ grows slowly with $N$. The variance of the average of $R$ RQMC samples using $n=N/R$ is
$\Theta(R^{-1}n^{-\theta})=\Theta(N^{-1}n^{1-\theta})=\Theta(N^{-1}N^{(1-\theta)/(\theta+1)})=\Theta(N^{-2\theta/(\theta+1)}).$
By requiring a guaranteed confidence level, rather than an asymptotic one, we get a higher variance for $\hat{\mu}$ than $\Theta(N^{-\theta})$ that we could have had with $R=O(1)$. For $\theta=2$ the variance rises from $\Theta(N^{-2})$ to $\Theta(N^{-4/3})$ and for $\theta=3$, it rises from $\Theta(N^{-3})$ to $\Theta(N^{-3/2})$. Put another way, obtaining a nonasymptotic confidence level raises the variance by $\Theta(N^{2/3})$ and $\Theta(N^{3/2})$ in those two cases.
While we cannot always know the best value of $n$ to use in a given setting, Theorem 1 gives us some guidance. The asymptotic value of $\theta$ may be known from theory. We may also have experience with similar integrands to make a judgment about what value of $\theta$ is reasonable in a given setting, or we can do some preliminary sampling to get a working value of $\theta$ before constructing a betting confidence interval. For any choice of $\theta$, that interval will be valid, though possibly wider than an interval using the true optimal $n$. Because $0\leqslant Y_{i}\leqslant 1$, we know that $\sigma_{0}^{2}\leqslant 1/4$. Therefore under the model (9), Theorem 1 provides the guidance to take
$n\leqslant\Bigl{(}\frac{9(\theta-1)^{2}N}{8\log(2/\alpha)}\Bigr{)}^{1/(\theta+1)}.$ (14)
For example, with $N=2^{10}$ and $\alpha=0.05$, the bound in (14) is always below $7$ whenever $1\leqslant\theta\leqslant 3$. The best integer value is $7$ for some values of $\theta$. For higher confidence levels (smaller $\alpha$), the optimal $n$ is smaller still.
The values of $N$ and $\alpha$ enter the Bennett half width of (13) only through the expression
$N_{\alpha}\equiv\frac{N}{\log(2/\alpha)}.$
Thus we can think of smaller $\alpha$ as providing a smaller effective budget $N_{\alpha}$.
Theorem 2.
Under the conditions of Theorem 1 with $\theta>1$ and $\sigma_{0}>0$, let $H_{\mathrm{MC}}$ be the half width $H^{\mathrm{Ben}}(1,N)$ using $n=1$ and $H_{\mathrm{RQMC}}$ be the half width $H^{\mathrm{Ben}}(n,N/n)$ using $n=n_{*}$ from (12). Let $N_{\alpha}=N/\log(2/\alpha)$. Then for $\theta>1$
$\frac{H_{\mathrm{RQMC}}}{H_{\mathrm{MC}}}$ $=\frac{\theta+1}{\sqrt{{2\sigma^{2}_{0}N_{\alpha}}}+1/3}\biggl{(}\frac{1}{2}[{3(\theta-1)}]^{1-\theta}N_{\alpha}\sigma_{0}^{2}\biggr{)}^{1/(\theta+1)}$
$=\Theta(N^{\frac{1-\theta}{2\theta+2}})$
as $N\to\infty$.
Proof.
Using equation (13) in the numerator, the ratio of these widths is
$\rho=\frac{H_{\mathrm{RQMC}}}{H_{\mathrm{MC}}}$ $=\frac{n^{(1-\theta)/2}\sqrt{2\sigma^{2}_{0}\log(2/\alpha)/N}+n\log(2/\alpha)/(3N)}{\sqrt{{2\sigma^{2}_{0}\log(2/\alpha)}/N}+\log(2/\alpha)/(3N)}$
$=\frac{n^{(1-\theta)/2}\sqrt{2\sigma^{2}_{0}N_{\alpha}}+n/3}{\sqrt{{2\sigma^{2}_{0}N_{\alpha}}}+1/3}.$
From equation (12)
$n^{(1-\theta)/2}$ $=({3(\theta-1)\sigma_{0}\sqrt{N_{\alpha}/2}})^{(1-\theta)/(\theta+1)}$
and we may also write $n=(3(\theta-1)\sigma_{0}(N_{\alpha}/2)^{1/2})^{2/(\theta+1)}$.
So for $\theta>1$, the width ratio is
$\rho$ $=\frac{({3(\theta-1)\sigma_{0}\sqrt{N_{\alpha}/2}})^{(1-\theta)/(\theta+1)}\sqrt{2\sigma^{2}_{0}N_{\alpha}}+(3(\theta-1)\sigma_{0}(N_{\alpha}/2)^{1/2})^{2/(\theta+1)}/3}{\sqrt{{2\sigma^{2}_{0}N_{\alpha}}}+1/3}$
$=\frac{({3(\theta-1)})^{(1-\theta)/(\theta+1)}2^{\theta/(\theta+1)}+(3(\theta-1))^{2/(\theta+1)}2^{-1/(\theta+1)}/3}{\sqrt{{2\sigma^{2}_{0}N_{\alpha}}}+1/3}(N_{\alpha}\sigma_{0}^{2})^{1/(\theta+1)}$
$=\frac{\theta+1}{\sqrt{{2\sigma^{2}_{0}N_{\alpha}}}+1/3}\biggl{(}\frac{1}{2}[{3(\theta-1)}]^{1-\theta}N_{\alpha}\sigma_{0}^{2}\biggr{)}^{1/(\theta+1)}$
as required. ∎
For the BVHK case with $\theta=2$, we find that RQMC narrows the widths by $\Theta(N^{-1/6})$. For the smooth case with $\theta=3$, we get a more favorable width ratio of $\Theta(N^{-1/4})$. The MC widths are $\Theta(N^{-1/2})$, and so for the BVHK case the RQMC widths are $\Theta(N^{-2/3})$ while for the smooth case, they are $\Theta(N^{-3/4})$.
4 Finite sample evaluations
Here we present some finite sample evaluations of the EBCI (specifically the predictable plug-in one) and the HBCI, both using RQMC, for 95% confidence. We begin with finite $N$ and for some integrands with known RQMC variance, we find the $n$ that minimizes the interval widths from Bennett’s inequality. They are then the exact best $n$ for Bennett’s inequality which the HBCIs and EBCIs of WauRam24a match asymptotically. After that we present finite sample results on the EBCIs and HBCIs for a variety of integrands differing in smoothness and dimension. The HBCIs are usually narrower. Finally, we compare the empirically optimal $n$ for the HBCIs to the values that optimize Bennett’s inequality. The HCBI widths approach those from Bennett’s inequality for large $R$ and we find that the optimal $n$ for HBCI tend to be the same or smaller than the ones we get from Bennett’s inequality.
For our HBCI computations, we used $\theta=1/2$. The software we used comes from hedged_algo which has code to reproduce the figures in WauRam24a . It calls the HBCI and predictable plug-in EBCI algorithms from bet_algo . The scrambled Sobol’ points we used were from QMCPy QMCPy2020a .
4.1 Semi-empirical evaluations
We can compute the width of intervals based on Bennett’s inequality for some specific cases. We saw in Section 3 that the RQMC sample sizes $n$ grow at slow rates $\Theta(N^{1/3})$ or $\Theta(N^{1/4})$ as the budget $N$ increases. Here we get a more detailed view of that phenomenon using the asymptotic formula for some specific variance rates. We can compute these confidence interval widths for very large values of $N$ that would be impractical to simulate.
The digital nets of Sobol’ Sob67 under the nested uniform scramble of Owe95 have a stratification property. The $n$ values of $x_{ij}$ for $i=1,\dots,n$ and any component $j=1,\dots,d$ have one value uniformly distributed in each interval $[(\ell-1)/n,\ell/n)$ for $\ell=1,\dots,n$ and those $n$ values are mutually independent. This allows us to work out the finite sample RQMC variances for some illustrative examples. The random linear scramble of Mat98 combined with a digital shift attains the same variance as the nested uniform scramble, but it has some dependencies among the stratified sample values. In both cases, the $d$ vectors $(x_{1j},\dots,x_{nj})\in\mathbb{R}^{n}$ are mutually independent.
We only consider values of $n$ that are powers of $2$ which is a best practice when using Sobol’ sequences Owe22a . The smooth error rate $\tilde{O}(n^{-3/2})$ for the RMSE is attainable along powers of $2$ but not for general $n$. An elegant argument due to Sobol’ sobo:1998 explains why. He notes that
$|f(\boldsymbol{x}_{n+1})-\mu|\leqslant(n+1)|\hat{\mu}_{n+1}-\mu|+n|\hat{\mu}_{n}-\mu|.$
A rate of $o(n^{-1})$ would imply that $f(\boldsymbol{x}_{n+1})$ itself consistently estimates $\mu$. For any sampling algorithm, that could only be true for a very minimal set of functions $f$. We also take $N$ to be a power of 2.
For $d=1$ and $f$ of bounded variation in the usual sense, we get $\operatorname{var}(\hat{\mu})\leqslant V_{\mathrm{HK}}(f)^{2}/n^{2}$. This holds because $D_{n}^{*}$ for the stratified sample cannot be larger than $1/n$. For illustration we choose a function with $V_{\mathrm{HK}}(f)=1$ for which we know exactly what the RQMC variance is for $n$ a power of $2$. That integrand is $f(x)=1\{x<1/3\}$. It is constant within $n-1$ of the intervals $[(\ell-1)/n,\ell/n)$. All of this function’s variation happens within just one of those intervals. By inspecting the base $2$ expansion $1/3=0.010101\cdots$ we can see that the remaining value of $f(x_{\ell})$ is $1$ or 00 and it equals $1$ with a probability that is either $1/3$ if $\log_{2}(n)$ is an odd integer or is $2/3$ if $\log_{2}(n)$ is an even integer. It follows that $\operatorname{var}(\hat{\mu}_{n})=2/(9n^{2})$.
Table 1 shows how the width minimizing value of $n$ increases with $N$ in the BVHK case with $f(x)=1\{x<1/3\}$. These are widths $W=2H^{\mathrm{Ben}}(n,N/n)$ using equation (11), not half widths. Some are larger than $1$, because the Bennett width does not reflect that we would intersect the intervals with $[0,1]$. The increase in $n$ is quite slow. By $N=1024$ the best RQMC sample size is only $n=8$ and this remains the optimal sample size up to $N=4096$. The value $N^{2/3}W$ very quickly becomes constant. For non-integer $\theta$ (not shown here) the scaled widths fluctuation within a finite range instead of approaching a constant.
The result in Table 1 for $N=1024$ is to choose $n=8$ which violates the guidance that the best $n$ is at most $7$ when $N=1024$ and $\alpha=0.05$. The reason is that this table restricts to powers of $2$, and $n=8$ is better than $n=4$.
$\log_{2}(N)$ $N$ $n$ $W$ $N^{2/3}W$
00 1 1 5.02e$+$00 5.020114
04 16 2 7.60e--01 4.826379
07 128 4 1.90e--01 4.826379
10 1,024 8 4.75e--02 4.826379
13 8,192 16 1.19e--02 4.826379
16 65,536 32 2.97e--03 4.826379
19 524,288 64 7.42e--04 4.826379
22 4,194,304 128 1.86e--04 4.826379
25 33,554,432 256 4.64e--05 4.826379
28 268,435,456 512 1.16e--05 4.826379
Table 1: As $N$ varies for $f(x)=1\{x<1/3\}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{2/3}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
Our second example is $f(x)=xe^{x-1}$. This function is bounded between 00 and $1$. We chose it because it is smooth enough to have $\theta=3$ while also not a polynomial or an antisymmetric function which would make the problem artificially easy for some methods. Because of the stratification, we know that
$\operatorname{var}(\hat{\mu})=\frac{1}{n^{2}}\sum_{\ell=1}^{n}\operatorname{var}(f(x_{\ell}))$ (15)
where $x_{\ell}\sim\mathbb{U}[(\ell-1)/n,\ell/n]$ are independent. Also
$\operatorname{var}(f(x_{\ell}))=n\int_{(\ell-1)/n}^{\ell/n}f(x)^{2}\,\mathrm{d}x-\Bigl{(}n\int_{(\ell-1)/n}^{\ell/n}f(x)\,\mathrm{d}x\Bigr{)}^{2}$ (16)
which can be computed for finite $n$ using the closed form antiderivatives of $f$ and $f^{2}$. We know that as $n$ increases through powers of 2
$\operatorname{var}(\hat{\mu})\asymp\frac{1}{12n^{3}}\int_{0}^{1}f^{\prime}(x)^{2}\,\mathrm{d}x=\frac{5-e^{-2}}{48n^{3}}$
but the exact variances are somewhat different for $n\leqslant 8$.
Table 2 shows how the optimal sample sizes $n$ grows with $N$ for our smooth example $f(x)=xe^{x-1}$. An RQMC sample size of $n=8$ becomes optimal for $N=4{,}096$ and remains optimal up to $N=32{,}768$. The values $N^{3/4}W$ settle down rapidly as $N$ increases. Consistent with the results of Section 3, the values $n$ grow at a slower rate with $N$ than in the BVHK case while the widths $W$ decrease at a faster rate.
$\log_{2}(N)$ $N$ $n$ $W$ $N^{3/4}W$
0 1 1 4.00e$+$00 4.003728
4 16 2 5.17e--01 4.138086
8 256 4 6.52e--02 4.175717
12 4,096 8 8.17e--03 4.185407
16 65,536 16 1.02e--03 4.187848
20 1,048,576 32 1.28e--04 4.188459
24 16,777,216 64 1.60e--05 4.188612
28 268,435,456 128 2.00e--06 4.188651
Table 2: As $N$ varies for $f(x)=xe^{x-1}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{3/4}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
4.2 Ridge functions
While the betting intervals are asymptotically of the same width as we would get from Bennett’s formula, here we investigate what choices of $n$ and $R$ give the narrowest interval widths for some finite $N$. These computations produce actual intervals that we then intersect with $[0,1]$. It is well known that smoothness and dimension affect the performance of RQMC methods. We choose to use ridge functions because that makes it straightforward to vary both the dimension and smoothness of our test functions.
Let $\Phi$ be the $\mathcal{N}(0,1)$ cumulative distribution function. Then for $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ the vector $\boldsymbol{z}=\Phi^{-1}(\boldsymbol{x})$ componentwise, has the $\mathcal{N}(0,I)$ distribution over $\mathbb{R}^{d}$. Then $\boldsymbol{1}^{\mathsf{\scriptsize T}}\boldsymbol{z}/\sqrt{d}\sim\mathcal{N}(0,1)$. Our ridge function
$f(\boldsymbol{x})=g\Bigl{(}\frac{\boldsymbol{1}^{\mathsf{\scriptsize T}}\Phi^{-1}(\boldsymbol{x})}{\sqrt{d}}\Bigr{)}.$
then has the same mean and variance and differentiability properties in any dimension $d$. We choose
$g_{\mathrm{jump}}(v)$ $=1\{v\geqslant 1\},$
$g_{\mathrm{kink}}(v)$ $=\frac{\min(\max(-2,v),1)+2}{3},$
$g_{\mathrm{smooth}}(v)$ $=\Phi(v+1),\quad\text{and}$
$g_{\mathrm{finance}}(v)$ $=\min\Bigl{(}1,\frac{\sqrt{\max(v+2,0)}}{2}\Bigr{)}.$
All of the above choices give ridge functions bounded between 00 and $1$. They span a very wide range of QMC regularity conditions. Using $g_{\mathrm{jump}}$ makes $f$ discontinuous and it has infinite variation in the sense of Hardy and Krause for $d\geqslant 2$. Using $g_{\mathrm{kink}}$ makes $f$ have infinite variation in the sense of Hardy and Krause for $d\geqslant 3$ and yet it has been seen to be a more ‘QMC friendly’ integrand than $g_{\mathrm{jump}}$ in high dimensions because it has an asymptotically bounded mean dimension mdridge , meaning that it is dominated by low dimensional interactions. It is continuous, has one weak derivative but is not differentiable everywhere. Using $g_{\mathrm{smooth}}$ makes $f$ infinitely differentiable which is generally a very QMC friendly property. Using $g_{\mathrm{finance}}$ gives $f$ a property similar to some integrands in financial option valuation, namely an unbounded derivative where $v=-2$.
For each of these four ridge functions, we considered budgets $N=2^{K}$ for $K\in\{8,10,12,14,16\}$ with RQMC sample sizes $n=2^{k}$ for $k\in\{0,1,2,3,4,5,6\}$ in dimensions $d\in\{1,2,4,16\}$. All of those conditions were repeated $20$ times independently. In each repetition we computed the width of the HBCI, the width of the EBCI, and the width of the CLT window.
The HBCI were narrower than the corresponding EBCI about 92.1% of the time. The CLT intervals were narrower than the corresponding HBCI 99.6% of the time. The few times that the CLT intervals were wider were predominantly from cases with $n=1$. The EBCI were never narrower than the CLT intervals. In some cases, the EBCI had width $1$. This only happened for $N=256$ and $n=32$ or $64$. Those cases have small replicate numbers $R=8$ and $R=4$, respectively.
The HBCI are most interesting because they have non-asymptotic coverage and are usually narrower than the EBCI. We made an extensive graphical exploration of their width versus $n$ (for every combination of dimension, ridge function and $N$). We similarly explored width versus budget $N$ and width versus dimension $d$, in each case at all levels of all of the other variables.
It was evident that the dimension $d$ made hardly any difference to the width of the confidence intervals. An additive model using factors equal to the dimension $d$, the RQMC sample size $n$, the function $f$ and the budget $N$ explained 94.09% of the variance in the logarithm of the HBCI width. That same model without dimension explained 94.05% of this variance. It was surprising to see no dimension effect for the HBCI width in our ridge functions. We knew that the mean and variance of $f(\boldsymbol{x})$ for any of those ridge functions does not depend on $d$ but that does not imply that averages over $n$ RQMC points would be invariant to $d$. Because $d$ made no material difference to the widths, we pooled information from all dimensions $d$.
It was clear graphically, and not surprising, that increasing $N$ by a factor of $4$ always narrows the HBCI width by a noticeable amount. The interesting non-monotone patterns were from the way that the width depends on $n$ with the other variables fixed. Figure 1 shows the mean widths of HBCI versus $n$ for all four ridge functions and all five budgets $N$.
Figure 1: Mean confidence interval width for the HBCI as a function of the RQMC sample size $n$. There is one panel for each of the $4$ ridge functions. Within a panel the curves are for $N=2^{2r}$ for $r=4,5,6,7,8$ from top to bottom. The smallest mean width for each $N$ is marked with a larger circle. The means are taken over all $20$ replicates and all $4$ dimensions used.
We see in Figure 1 some things that align very well with our analysis and some that were not anticipated. The optimal samples sizes $n$ are very small and they grow slowly with the budget $N$ as expected. The least smooth ridge function $g_{\mathrm{jump}}$ has the fastest growing $n$ as expected. The interval widths for $g_{\mathrm{finance}}$ and $g_{\mathrm{smooth}}$ were very close to each other despite having very different smoothness and then they had identical optimal $n$ at each $N$. For $N=2^{10}$ the optimal $n$ for $g_{\mathrm{kink}}$ is larger than the one for $g_{\mathrm{smooth}}$ as expected but at $N=2^{16}$, it is $g_{\mathrm{smooth}}$ that has the larger optimal $n$.
Our expectations were based on asymptotic considerations. First, we used the fact that the HBCI widths asymptotically approach the optimal EBCI widths from Bennett’s inequality. Second, we supposed that the RQMC variance is asymptotic to $\sigma_{0}^{2}n^{-\theta}$ apart from logarithmic factors. We also used the finite sample Bennett widths from (13) because the asymptotic expression $\sigma_{n}\sqrt{2\log(2/\alpha)/R}$ does not have a meaningful optimum. The empirical results strongly confirm our asymptotic findings that small $n$ is best overall, but among such small $n$, it is less clear how to choose $n$ based on smoothness of $f$.
4.3 Functions with known RQMC variance
Here we return to the functions from Section 4.1 with known non-asymptotic RQMC variance. They allow us to directly compare the widths of the HBCI to the value we get from Bennett’s formula. We know that the HBCI have lengths that asymptotically match Bennett’s formula, but we also want to study how those lengths approach Bennett’s formula in some examples.
Using the known variances, we can compute the half width from Bennett’s inequality  (11) for our $\sigma^{2}_{n}$ and $R=N/n$ double it and then divide the average HBCI width of 20 independent trials by this value. The results are shown in Figure 2. For our two example functions we see that as $R=N/n$ increases (so $n$ decreases for fixed $N$) the ratio of widths decreases and approaches the theoretically expected value of $1$. At the smallest $R$ (largest $n$) the HBCI were two or three times as wide as the Bennett intervals for the discontinuous function. They were somewhat wider for the smooth function.
Figure 2: Ratio of mean HBCI width to width from Bennett’s inequality as a function of the RQMC sample size $n$ for the given budgets $N$. The means for the HBCI are taken over the 20 independent replicates.
The HBCI are computed without knowledge of the exact variance $\sigma^{2}_{n}$ that the Bennett interval uses. They are instead based on estimates of that variance and they must therefore allow for uncertainty in the variance estimate. It is then not surprising that the resulting intervals are wider than the Bennett ones. We see in Table 3 that the HBCI’s minimum widths were not found at larger $n$ than the minimizers of the Bennett intervals. They were either equal to or half as large as $n_{*}$. We also see that empirical best values of $n$ are no larger for the smooth function than for the discontinuous one, as expected.
$1\{x<1/3\}$ $xe^{x-1}$
$N$ $n_{*}$ $n_{\mathrm{opt}}$ $n_{*}$ $n_{\mathrm{opt}}$
256 4 4 4 2
1,024 8 4 4 4
4,096 8 8 8 4
16,384 16 16 8 8
65,536 32 16 16 8
Table 3: For the given budgets $N$, we show $n_{*}$ as the minimizer of the Bennett width (13) over $n=2^{k}$ for integer $k$ and $n_{\operatorname{opt}}$ is the value of $n$ that minimized the average width of the HBCI for the given $N$. These are shown for functions $f(x)=1\{x<1/3\}$ and $f(x)=xe^{x-1}$.
5 Discussion
In favorable cases the RQMC variance is $\tilde{O}(n^{-3})$ and then using a fixed number $R$ of replicates of point sets with $n\to\infty$ we can estimate the integral $\mu$ with a standard deviation of $\tilde{O}(N^{-3/2})$ for $N=nR$ along with an unbiased estimate of the variance of that estimate. It remains very difficult to get a confidence interval of width $\tilde{O}(N^{-3/2})$ err4qmc . Using Student’s $t$ with a small $R$ gave good results in an extensive simulation by LEcEtal24a but that success is not yet theoretically understood. For an integrand subject to known bounds, we can get a non-asymptotic confidence interval by using $R$ replicates of $n$ RQMC points using a predictable plug-in empirical Berstein confidence interval. There, if $\sigma^{2}_{n}=\Theta(n^{-\theta})$ then the optimal $n$ is $\Theta(N^{1/(\theta+1)})$ and the resulting confidence intervals have width $\Theta(N^{-\theta/(\theta+1)})$. While it remains to find a way to choose the optimal $n$ empirically, we have strong theoretical guidance from equation (14) that we can interpret as an upper bound on $n$ for the oracle which is then a reasonable upper bound for HBCIs which have to estimate the RQMC variance. When $n$ must be a small power of $2$, we can rule out many suboptimal choices.
We have noted several times that there is an asymptotic equivalence between the Bennett interval widths and those of the EBCI and HBCI intervals. In our setting that equivalence is as $R\to\infty$ for fixed $n$. In the RQMC context, the best $n$ is quite small. Then for large $N$ it is not surprising that an asymptote as $R\to\infty$ is very predictive of our results.
It also remains to find a good way to use RQMC in confidence sequences as opposed to confidence intervals. We might set up $R$ independent infinite sequences of RQMC points and then stop them at the first $n$ where they provide a $1-\alpha$ interval narrower than some $\epsilon$. In a task like that it makes sense to only use $n=2^{k}$ for integers $k$. That is partly because powers of two are good for scrambled Sobol’ points, but also because in RQMC it is generally advisable to study sample sizes $n$ that grow geometrically not arithmetically sobo:1998 . The idea is that if $n$ is not large enough to get a good answer, then $n+1$ is unlikely to be much of an improvement. A confidence sequence based on geometrically growing $n$ would not have to be as wide as one for arithmetically growing $n$.
For scrambled Sobol’ points it is strongly advisable to take $n=2^{k}$ because those are the best sample sizes and they can even give a better convergence rate. The Halton sequence Hal60 does not have strongly superior sample sizes and so $n$ need not be a power of two for it. The two scrambles we considered for Sobol’ points can also be applied to the Halton sequence getting $\operatorname{var}(\hat{\mu})=o(1/n)$ without requiring special sample sizes $n$ haltongain . There are however no published non-trivial settings where scrambled Halton points can obtain $\theta>2$.
Acknowledgments
We thank Aaditya Ramdas and Ian Waudby-Smith for helpful discussions. ABO was supported by the U.S. National Science Foundation grant DMS-2152780. FJH and AGS were supported by U.S. National Science Foundation grant DMS-2316011. AGS was supported by the U.S. DOE Office of Science Graduate Student Research Program.

### v2

@@
 We study the combination of randomized quasi-Monte Carlo (RQMC) integration to estimate an expectation with some non-asymptotic (finite sample valid) methods to get a confidence interval for such an expectation.
-RQMC has significant accuracy benefits over plain Monte Carlo (that we outline below), but the usual confidence intervals based on RQMC estimates have justifications that are asymptotic in their sample size LEcEtal24a .
+RQMC has significant accuracy benefits over plain Monte Carlo (that we outline below), but the usual confidence intervals based on RQMC estimates have justifications that are asymptotic in their sample size Nakayama and Tuffin (2024).
 Empirical Bernstein and some related betting methods (described below) give finite sample assurances for bounded integrands.
-Specifically, if it is known that the integrand values must lie between 00 and $1$, then those methods can construct a confidence interval that has at least $1-\alpha$ probability to contain the integral’s value.
+Specifically, if it is known that the integrand values must lie between 0 and $1$, then those methods can construct a confidence interval that has at least $1-\alpha$ probability to contain the integral’s value.
 This confidence level holds for any integrand satisfying the given bounds.
-The context for our work is as follows.
-Many scientific problems require the numerical evaluation of multidimensional integrals.
-When the dimension is high enough then classical tensor product integration rules like those in DavRab84 become too expensive to use.
-It is then common to use plain Monte Carlo (MC) methods instead.
-MC methods based on random sampling typically have root mean squared errors (RMSEs) of $O(n^{-1/2})$ given $n$ function evaluations.
-This rate is slow but it is the same in any dimension.
-MC also allows one to use statistical methods to quantify the uncertainty in the estimated integral.
-Quasi-Monte Carlo (QMC) methods (e.g., DicPil10a and Nie92 ) are deterministic sampling strategies that under a bounded variation assumption produce integral estimates with error $\tilde{O}(n^{-1})$ where $\tilde{O}(\cdot)$ means we neglect powers of $\log(n)$.
-Plain QMC methods are deterministic and so they lose the uncertainty quantification advantage of MC.
-RQMC methods, surveyed in LEcLem02a , allow one to make independent replicates of a statistically unbiased QMC method to support variance estimates and asymptotic confidence intervals.
-RQMC methods can also improve on the convergence rate of QMC methods obtaining an RMSE of $\tilde{O}(n^{-3/2})$ under a smoothness assumption Owe97 ; Owe08a on the integrand.
+[…]
 In addition to asymptotic confidence interval methods for RQMC, some other papers have provided finite sample uncertainty quantifications based on additional assumptions.
-For purely independent and identically distributed (IID) sampling, knowing a bound on the kurtosis facilitates a finite sample confidence interval for the mean of a random variable in terms of the sample mean and the sample standard deviation HicEtal14a .
-For deterministic QMC sampling of $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ with lattices or digital nets, assuming that the coefficients in a suitable orthogonal decomposition of $f:[0,1]^{d}\to\mathbb{R}$ decay reasonably, one may derive a deterministic error bound on $\mu=\mathbb{E}[f(\boldsymbol{x})]$ HicJim16a ; HicEtal17a ; JimHic16a ; JagSor23a .
+For purely independent and identically distributed (IID) sampling, knowing a bound on the kurtosis facilitates a finite sample confidence interval for the mean of a random variable in terms of the sample mean and the sample standard deviation Hickernell et al.
+(2013).
+For deterministic QMC sampling of $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ with lattices or digital nets, assuming that the coefficients in a suitable orthogonal decomposition of $f:[0,1]^{d}\to\mathbb{R}$ decay reasonably, one may derive a deterministic error bound on $\mu=\mathbb{E}[f(\boldsymbol{x})]$ Hickernell and Jiménez Rugama (2016); Hickernell et al.
+(2018); Jiménez Rugama and Hickernell (2016); Sorokin and Jagadeeswaran (2024).
 A different approach is to assume that $f$ is a realization of a Gaussian process and construct a credible interval for $\mathbb{E}[f(\boldsymbol{x})]$.
-If the covariance kernel and the sampling nodes are well-matched, the computational effort required is nearly $O(n)$ and not the $O(n^{3})$ effort typically required for such credible intervals RatHic19a ; JagHic22a .
-The GAIL Gail_ug and QMCPy QMCPy2020a libraries implement these finite sample uncertainty quantifications under the assumptions outlined above.
+If the covariance kernel and the sampling nodes are well-matched, the computational effort required is nearly $O(n)$ and not the $O(n^{3})$ effort typically required for such credible intervals Jagadeeswaran and Hickernell (2019, 2022).
+The GAIL Choi et al.
+(2015) and QMCPy Choi et al.
+(2025) libraries implement these finite sample uncertainty quantifications under the assumptions outlined above.
 Here, we study RQMC confidence intervals assuming that the integrand is bounded between zero and one.
 The advantage of this assumption is that there are more settings where we are certain that it holds.
-The non-asymptotic confidence intervals that we emphasize are the recently developed hedged betting confidence intervals (HBCI) and some related predictable plug-in empirical Bernstein confidence intervals (EBCI), both from WauRam24a .
+The non-asymptotic confidence intervals that we emphasize are the recently developed hedged betting confidence intervals (HBCI) and some related predictable plug-in empirical Bernstein confidence intervals (EBCI), both from Waubdy-Smith and Ramdas (2024).
 They are derived from some infinite confidence sequences and we believe that they provide the narrowest confidence intervals among currently available methods.
-Section 2 gives our notation and some definitions and background on quasi-Monte Carlo, randomized quasi-Monte Carlo, and the EBCI and HBCI of WauRam24a .
+Section 2 gives our notation and some definitions and background on quasi-Monte Carlo, randomized quasi-Monte Carlo, and the EBCI and HBCI of Waubdy-Smith and Ramdas (2024).
 In Section 3 we show how splitting $N$ observations into $R=N/n$ independent replicates of $n$ RQMC point sets can give narrower EBCI than MC does for an oracle that knows the RQMC variance at each $n$.
@@
 Section 4 makes some computational investigations.
-As predicted by the theory the best values of $n$ to use grew slowly with $N$.
-The empirically best values of $n$ for HBCI were no larger than the oracle values, and sometimes smaller.
+As predicted by the theory the best values of $n$ to use grow slowly with $N$.
+The empirically best values of $n$ for HBCI are no larger than the oracle values, and sometimes smaller.
 Section 5 has some final comments and discussion.
@@
 We use $1\{E\}$ to denote a quantity that equals $1$ when expression $E$ holds and is zero otherwise.
-The vectors $\boldsymbol{0}$ and $\boldsymbol{1}$ have all components equal to 00 and $1$ respectively with a dimension given by context.
+The vectors $\boldsymbol{0}$ and $\boldsymbol{1}$ have all components equal to 0 and $1$ respectively with a dimension given by context.
 We write $a_{n}=O(b_{n})$ as $n\to\infty$ if there exists $C>0$ and $n_{0}<\infty$ with $|a_{n}|\leqslant Cb_{n}$ whenever $n\geqslant n_{0}$.
@@
 $\Pr(A\leqslant\mu\leqslant B)\geqslant 1-\alpha.$ (1)
-A confidence interval is strict when the right hand side above is an equality. An asymptotic confidence interval is a sequence $(A_{n},B_{n})$ of random quantities for which
+A confidence interval is strict when the sides of (1) are equal. An asymptotic confidence interval is a sequence $(A_{n},B_{n})$ of random quantities for which
 $\lim_{n\to\infty}\Pr(A_{n}\leqslant\mu\leqslant B_{n})\geqslant 1-\alpha.$
@@
 $\Pr\biggl(\sup_{n\geqslant 1}S_{n}\geqslant\eta\biggr)\leqslant\frac{\mathbb{E}(S_{1})}{\eta}.$
-The confidence sequences from WauRam24a are derived by applying Ville’s inequality vill:1939 to nonnegative supermartingales.
+The confidence sequences from Waubdy-Smith and Ramdas (2024) are derived by applying Ville’s inequality Ville (1939) to nonnegative supermartingales.
 The problem we consider is to approximate the finite-dimensional integral
@@
 $\lim_{n\to\infty}\Pr\Bigl(\sqrt{n}\frac{\hat{\mu}-\mu}{s}\leqslant t^{1-\alpha}_{(n-1)}\Bigr)=1-\alpha$
-where $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on $n-1$ degrees of freedom.
+where $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on n1n-1 degrees of freedom.
 Then $\hat{\mu}\pm st^{1-\alpha/2}/\sqrt{n}$ is an asymptotic $1-\alpha$ confidence interval for $\mu$.
@@
 $D_{n}^{*}=D_{n}^{*}(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})=\sup_{\boldsymbol{a}\in[0,1]^{d}}\,\biggl|\frac{1}{n}\sum_{i=1}^{n}1\{\boldsymbol{x}_{i}\in[\boldsymbol{0},\boldsymbol{a})\}-\prod_{j=1}^{d}a_{j}\biggr|.$
-Small values of $D_{n}^{*}$ show that all anchored boxes $[\boldsymbol{0},\boldsymbol{a})$ have very nearly the desired proportion of the $n$ points, which is the volume of $[\boldsymbol{0},\boldsymbol{a})$.
-It is possible to choose points $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}$ so that $D_{n}^{*}=\tilde{O}(n^{-1})$ while plain MC points have $D_{n}^{*}=\tilde{O}(n^{-1/2})$ Nie92 .
-The improvement in star discrepancy yields an improvement in integration error via the Koksma-Hlawka inequality Hic04a :
+Small values of $D_{n}^{*}$ show that all anchored boxes $\boldsymbol{0},\boldsymbol{a})$ have very nearly the desired proportion of the $n$ points, which is the volume of $\boldsymbol{0},\boldsymbol{a})$.
+It is possible to choose points $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}$ so that $D_{n}^{*}=\tilde{O}(n^{-1})$ while plain MC points have $D_{n}^{*}=\tilde{O}(n^{-1/2})$ Niederreiter (1992).
+The improvement in star discrepancy yields an improvement in integration error via the Koksma-Hlawka inequality Hickernell (2006):
 $|\hat{\mu}-\mu|\leqslant D_{n}^{*}(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})\times V_{\mathrm{HK}}(f).$ (3)
@@
 The new factor $V_{\mathrm{HK}}(f)$ is the total variation of $f$ in the sense of Hardy and Krause.
-See Owe05a .
+See Owen (2005).
 It follows from (3) that we can get $|\hat{\mu}-\mu|=\tilde{O}(n^{-1})$ using QMC.
@@
 $\hat{\mu}=\frac{1}{R}\sum_{i=1}^{R}\hat{\mu}_{i}\quad\text{and}\quad S^{2}=\frac{1}{R-1}\sum_{i=1}^{R}(\hat{\mu}_{i}-\hat{\mu})^{2}.$
-Here $S^{2}/R$ is an unbiased estimate of the MSE of the pooled estimate $\hat{\mu}$. The confidence level $1-\alpha$ is obtained in the limit as $R\to\infty$ naka:tuff:2024 for fixed $n$.
+Here $S^{2}/R$ is an unbiased estimate of the MSE of the pooled estimate $\hat{\mu}$.
+The confidence level $1-\alpha$ is obtained in the limit as $R\to\infty$ Nakayama and Tuffin (2024) for fixed $n$.
+That convergence is generally quicker for more nearly Gaussian $\hat{\mu}_{i}$.
 In addition to getting an unbiased estimate of the RQMC variance, randomization can provide some other benefits.
 First, some RQMC methods can attain better accuracy than plain QMC.
-Under sufficient smoothness, the scrambled digital sequences in Owe95 attain RMSEs of $\tilde{O}(n^{-3/2})$ Owe97 ; Owe08a .
-The same holds for the scramblings of Mat98 .
+Under sufficient smoothness, the scrambled digital sequences in Owen (1995) attain RMSEs of $\tilde{O}(n^{-3/2})$ Owen (1997, 2008).
+The same holds for the scramblings of Matoušek (1998).
 We will refer to this as the ‘smooth case’.
-Higher order digital nets Dic11a can attain even better convergence rates, though the rates are not necessarily evident empirically when $d$ is moderately large, which nuyens2010higher attribute to numerical precision.
+Higher order digital nets Dick (2011) can attain even better convergence rates, though the rates are not necessarily evident empirically when $d$ is moderately large, which Nuyens and Cools (2010) attribute to numerical precision.
 Those RQMC methods have further useful properties.
@@
 Furthermore, the condition of finite variation in the sense of Hardy and Krause can be quite strict.
-For example, when $d\geqslant 2$, step discontinuities in the integrand $f$ ordinarily make $V_{\mathrm{HK}}(f)$ infinite unless those discontinuities are axis parallel Owe05a .
+For example, when $d\geqslant 2$, step discontinuities in the integrand $f$ ordinarily make $V_{\mathrm{HK}}(f)$ infinite unless those discontinuities are axis parallel Owen (2005).
 This makes QMC unattractive for estimating integrals of binary valued functions, while RQMC will still attain an RMSE that is $o(n^{-1/2})$.
-While there are numerous RQMC methods in use, we will restrict ourselves here to scramblings of digital nets and sequences, such as the ones of Sobol’ Sob67 .
-Our variance formulas assume that the scramblings are either those of Owe95 or Mat98 .
-In our computations we used the linear matrix scramble of Mat98 with a digital shift.
-While RQMC methods yield very good accuracy, most of the known confidence interval methods for them supply only asymptotic confidence as $R\to\infty$ naka:tuff:2024 .
-In some very limited settings there are central limit theorems for $\hat{\mu}$ as $n\to\infty$ Loh01 .
+While there are numerous RQMC methods in use, we will restrict ourselves here to scramblings of digital nets and sequences, such as the ones of Sobol’ Sobol’ (1967).
+Our variance formulas assume that the scramblings are either those of Owen (1995) or Matoušek (1998).
+In our computations we used the linear matrix scramble of Matoušek (1998) with a digital shift.
+While RQMC methods yield very good accuracy, most of the known confidence interval methods for them supply only asymptotic confidence as $R\to\infty$ Nakayama and Tuffin (2024).
+In some very limited settings there are central limit theorems for $\hat{\mu}$ as $n\to\infty$ Loh (2003).
 Having the confidence statements be asymptotic in $R$ is unfortunate since the RMSE of $\hat{\mu}$ is proportional to $R^{-1/2}$ while the RMSE vanishes at a faster rate in $n$.
@@
 It is not desirable to make such assumptions and so we prefer a nonparametric method with non-asymptotic confidence.
-There is a theorem of Bahadur and Savage BahSav56 that reveals sharp constraints on our ability to construct nonparametric and non-asymptotic confidence intervals.
-Here is the description from err4qmc :
-They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)$ be $\mathbb{E}(Y)$ when $Y\sim F\in\mathbb{R}$, their conditions are:
+There is a theorem of Bahadur and Savage Bahadur and Savage (1956) that reveals sharp constraints on our ability to construct nonparametric and non-asymptotic confidence intervals.
+Here is the description from Owen (2026):
+They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)=\mathbb{E}(Y)$ for $Y\sim F\in\mathcal{F}$, their conditions are:
   1.  (i)
@@
   3.  (iii)
-    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $10<\pi<$ 1, then $\pi F+(1-\pi)G\in\mathcal{F}$.
+    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $0<\pi<1$, then $\pi F+(1-\pi)G\in\mathcal{F}$.
 Then their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in\mathbb{R}$ with probability at least $1-\alpha$.
@@
 $\bar{Y}_{R}\pm\sqrt{\frac{\log(2/\alpha)}{2R}}$ (4)
-for $10<\alpha<$ 1 is a $1-\alpha$ confidence interval for $\mu=\mathbb{E}(Y_{i})$.
-Equation (4) gives the confidence interval of Hoeffding hoef:1963 .
+for $0<\alpha<1$ is a $1-\alpha$ confidence interval for $\mu=\mathbb{E}(Y_{i})$.
+Equation (4) gives the confidence interval of Hoeffding Hoeffding (1963).
 The interval is not always nested within $[0,1]$, but we can simply take its intersection with $[0,1]$.
@@
 One approach is to make use of $\sigma^{2}=\operatorname{var}(Y)$ if it is known.
-Hoeffding’s confidence interval is a consequence of Hoeffding’s exponential probability bound in hoef:1963 .
+Hoeffding’s confidence interval is a consequence of Hoeffding’s exponential probability bound in Hoeffding (1963).
 Other exponential probability bounds take advantage of a known value $\sigma^{2}=\operatorname{var}(Y)$ or even a sample estimate of that variance.
 Those bounds can be used to get improved confidence intervals.
-Theorem 3 of Maurer and Pontil maurer2009empirical is that
+Theorem 3 of Maurer and Pontil Maurer and Pontil (2009) is that
 $\mathbb{E}(Y)-\frac{1}{R}\sum_{i=1}^{R}Y_{i}\leqslant\sqrt{\frac{2\sigma^{2}\log(1/\delta)}{R}}+\frac{\log(1/\delta)}{3R}.$
 holds with probability at least $1-\delta$.
-They call this Bennett’s inequality, because it derives from an inequality in benn:1962 .
-They do not supply a proof, but it can be proved from Bernstein’s inequality given as Theorem 3 of bouc:lugo:bous:2003 .
+They call this Bennett’s inequality, because it derives from an inequality in Bennett (1962).
+They do not supply a proof, but it can be proved from Bernstein’s inequality given as Theorem 3 of Boucheron et al.
+(2003).
 Taking $\delta=\alpha/2$ we get
@@
 Replacing $\sigma^{2}$ by $s^{2}$ in (5) would not always give a valid confidence interval.
-Theorem 4 of maurer2009empirical shows that for $R\geqslant 2$
+Theorem 4 of Maurer and Pontil (2009) shows that for $R\geqslant 2$
 $\bar{Y}_{R}\pm\sqrt{\frac{2s^{2}\log(4/\alpha)}{R}}+\frac{7\log(4/\alpha)}{3(R-1)}.$ (6)
@@
 We see that both terms on the right hand side of (6) are larger than the ones in the Bennett interval (5).
-Their interval has improved asymptotic constants compared to the original empirical Bernstein inequality from audi:muno:szep:2007 .
-The EBCI and HBCI from WauRam24a provide a further asymptotic improvement.
+Their interval has improved asymptotic constants compared to the original empirical Bernstein inequality from Audibert et al.
+(2007).
+The EBCI and HBCI from Waubdy-Smith and Ramdas (2024) provide a further asymptotic improvement.
 2.4 Betting methods
-Our interest in this problem was sparked in part by some recent work by WauRam24a on confidence intervals generated by betting arguments.
+Our interest in this problem was sparked in part by some recent work by Waubdy-Smith and Ramdas (2024) on confidence intervals generated by betting arguments.
 They consider a more general setting of confidence sequences, as in equation (2), that can then be specialized to confidence intervals as needed.
@@
 Then if $m\neq\mu$ somebody starting with a stake of \$1.00 can make a series of bets against $H_{0}(m)$ as the $Y_{i}$ are revealed in order and have an expected fortune that grows without bound.
-Just prior to time $i$, the bettor picks a value $\lambda_{i}\in(-1/(1-m),1/m)$.
+Just prior to time $i$, the bettor picks a value $\lambda_{i}=\lambda_{i}(m)\in(-1/(1-m),1/m)$.
 The bettor’s capital at time $t\geqslant 1$ is
@@
 To take $\lambda_{i}(m)<0$ is to bet that $Y_{i}<m$.
-Section 4 of WauRam24a notes that $\mathcal{K}_{t}(\mu)$ is a nonnegative martingale, so that $\mathbb{E}(\mathcal{K}_{t}(\mu))=1$ for all $t\geqslant 1$.
+Section 4 of Waubdy-Smith and Ramdas (2024) notes that $\mathcal{K}_{t}(\mu)$ is a nonnegative martingale, so that $\mathbb{E}(\mathcal{K}_{t}(\mu))=1$ for all $t\geqslant 1$.
 There is no way for the bettor to pick bet sizes $\lambda_{i}(\mu)$ to make an expected profit.
@@
 As a result we can get a confidence sequence by retaining at time $t$, all the values $m$ for which $\max_{1\leqslant\tau\leqslant t}\mathcal{K}_{\tau}(m)\leqslant 1/\alpha$.
-This idiom uses a hypothetically infinite ensemble of bettors to do this, but WauRam24a give implementable algorithms for it.
+This idiom uses a hypothetically infinite ensemble of bettors to do this, but Waubdy-Smith and Ramdas (2024) give implementable algorithms for it.
 The other side of the coin is that there are good betting strategies when $m\neq\mu$.
@@
 $\lambda_{t}^{\text{PrPl-EB}}=\sqrt{\frac{2\log(2/\alpha)}{\hat{\sigma}^{2}_{t-1}t\log(1+t)}}\,\wedge c$
-for some $c\in(0,1)$ with $c=1/2$ or $3/4$ given as reasonable defaults. Note that this method only bets that $Y_{i}>m$. Theorem 2 of WauRam24a gives a $1-\alpha$ confidence sequence
+for some $c\in(0,1)$ with $c=1/2$ or $3/4$ given as reasonable defaults.
+Note that this method only bets that $Y_{i}>m$.
+Theorem 2 of Waubdy-Smith and Ramdas (2024) gives a $1-\alpha$ confidence sequence
 $C_{t}^{\text{PrPl-EB}}=\frac{\sum_{i=1}^{t}\lambda_{i}Y_{i}}{\sum_{i=1}^{t}\lambda_{i}}\pm\frac{\log(2/\alpha)+\sum_{i=1}^{t}\nu_{i}\psi_{\mathrm{E}}(\lambda_{i})}{\sum_{i=1}^{t}\lambda_{i}},$ (7)
@@
 The running intersection $\cap_{1\leqslant\tau\leqslant t}C^{\text{PrPl-EB}}_{\tau}$ is also a $1-\alpha$ confidence sequence for $\mu$ and of course we can intersect any of these intervals with $[0,1]$.
-These confidence sequences can be specialized to confidence intervals when we have a fixed target sample size $R$ in mind. For that case WauRam24a recommends
+These confidence sequences can be specialized to confidence intervals when we have a fixed target sample size $R$ in mind. For that case Waubdy-Smith and Ramdas (2024) recommends
 $\lambda_{i}^{\text{PrPl-EB}(R)}=\sqrt{\frac{2\log(2/\alpha)}{R\hat{\sigma}^{2}_{i-1}}}\,\wedge c,\quad i=1,\dots,R.$ (8)
 The above formula for $\lambda_{i}$ uses the ordering of the data values and puts unequal weight on the $R$ different $Y_{i}$.
-There is a permutation strategy in WauRam24a to treat the $Y_{i}$ values more symmetrically, but they report that it makes little difference.
+There is a permutation strategy in Waubdy-Smith and Ramdas (2024) to treat the $Y_{i}$ values more symmetrically, but they report that it makes little difference.
 The EBCI we study take the form
 $\bigcap_{1\leqslant\tau\leqslant t}C_{t}^{\text{PrPl-EB}(R)}$
-where $C_{t}^{\text{PrPl-EB}(R)}$ is the interval in (7) using $\lambda_{i}$ from (8). Equation (17) of WauRam24a gives the scaled half width of these EBCI as
+where $C_{t}^{\text{PrPl-EB}(R)}$ is the interval in (7) using $\lambda_{i}$ from (8). Equation (17) of Waubdy-Smith and Ramdas (2024) gives the scaled half width of these EBCI as
 $\sqrt{R}\biggl(\frac{\log(2/\alpha)+\sum_{i=1}^{R}\nu_{i}\psi_{\mathrm{E}}(\lambda_{i})}{\sum_{i=1}^{R}\lambda_{i}}\biggr)\stackrel{{\scriptstyle\mathrm{a.s.}}}{{\to}}\sigma\sqrt{2\log(2/\alpha)}$
-as $R\to\infty$. The corresponding limit for the empirical Bernstein CIs of maurer2009empirical is $\sigma\sqrt{2\log(4/\alpha)}$.
-The hedged betting process of WacMenSch08 works as follows. For positive predictable sequences $\lambda_{i}^{+}(m)$ and $\lambda_{i}^{-}(m)$ defined below, they use capital processes
-$K_{t}^{+}(m)=\prod_{i=1}^{t}(1+\lambda_{i}^{+}(m)(Y_{i}-m)\quad\text{and}\quad K_{t}^{-}(m)=\prod_{i=1}^{t}(1-\lambda_{i}^{+}(m)(Y_{i}-m)$
+as $R\to\infty$. The corresponding limit for the empirical Bernstein CIs of Maurer and Pontil (2009) is $\sigma\sqrt{2\log(4/\alpha)}$.
+The hedged betting process of Wackerly et al. (2008) works as follows. For positive predictable sequences $\lambda_{i}^{+}(m)$ and $\lambda_{i}^{-}(m)$ defined below, they use capital processes
+$K_{t}^{+}(m)=\prod_{i=1}^{t}\bigl(1+\lambda_{i}^{+}(m)(Y_{i}-m)\bigr)\quad\text{and}\quad K_{t}^{-}(m)=\prod_{i=1}^{t}\bigl(1-\lambda_{i}^{+}(m)(Y_{i}-m)\bigr)$
 that bet, respectively, on $Y_{i}>m$ and $Y_{i}<m$. The hedged process is
@@
 For any $\theta$, $K_{t}^{\pm}(\mu)$ is a nonnegative martingale hence subject to Ville’s theorem.
-The supplement to WauRam24a suggests $\theta=1/2$ as a default.
+The supplement to Waubdy-Smith and Ramdas (2024) suggests $\theta=1/2$ as a default.
 If $m\neq\mu$ then $K_{t}^{\pm}(m)$ will tend to grow with $t$.
@@
 3 Asymptotic confidence interval comparisons
-Our confidence intervals will use $R$ IID replicates that each use $n$ RQMC points in $[0,1]^{d}$.
+Our confidence intervals will use $R$ IID replicates that each use $n$ RQMC points in $0,1]^{d}$.
 That takes $N=nR$ evaluations of $f$.
@@
 For the smooth case, $\sigma^{2}_{n}=\tilde{O}(n^{-3})$.
-We know from WauRam24a that the widths of their predictable plug-in EBCI approaches that of some Bennett intervals described below, as $R\to\infty$ for fixed $n$.
+We know from Waubdy-Smith and Ramdas (2024) that the widths of their predictable plug-in EBCI approaches that of some Bennett intervals described below, as $R\to\infty$ for fixed $n$.
 The same is true of their HBCI.
@@
 As mentioned above, the tilde in $\tilde{O}$ hides powers of $\log(n)$.
-Those powers are present in error bounds for adversarially chosen integrands but do not seem to come up in real problems schl:2002 ; wherearethelogs .
+Those powers are present in error bounds for adversarially chosen integrands but do not seem to come up in real problems Schlier (2002); Owen and Pan (2022).
 We will consider working models of the form
@@
 An oracle that knew $\sigma^{2}_{n}$ could use this formula to select the best $n$ and $R$ for a confidence interval subject to a constraint $nR=N$.
-The predictable plug-in EBCI and the HBCI from WauRam24a do not assume a known variance for the $Y_{i}$.
+The predictable plug-in EBCI and the HBCI from Waubdy-Smith and Ramdas (2024) do not assume a known variance for the $Y_{i}$.
 However those intervals attain the same asymptotic limiting width as $R\to\infty$ that an oracle would get from Bennett’s formula.
@@
 $H(n)=H^{\mathrm{Ben}}(n,N/n)=\sigma_{0}n^{(1-\theta)/2}N^{-1/2}\sqrt{2\log(2/\alpha)}+\frac{\log(2/\alpha)}{3}\frac{n}{N}.$ (13)
-If $\theta=1$, then $H(n)$ is minimized over $[1,\infty)$ at $n=1$.
+If $\theta=1$, then $H(n)$ is minimized over $1,\infty)$ at $n=1$.
 For $\theta>1$, the function $H(n)$ is a convex function of $n$. It has a unique minimum over $(0,\infty)$ at some $n_{*}$ where $H^{\prime}(n_{*})=0$. Now
@@
 If $1<n_{*}<N$ and $n_{*}$ is not an integer then the best integer $n$ is either $\lfloor n_{*}\rfloor$ or $\lceil n_{*}\rceil$ by the convexity noted in the proof.
-For scrambled Sobol’ points it is generally best to take $n$ to be a power of 2 Owe22a , which is not reflected in Theorem 1. We incorporate that restriction into some numerical examples in Section 4.
+For scrambled Sobol’ points it is generally best to take $n$ to be a power of 2 Owen (2022), which is not reflected in Theorem 1.
+We incorporate that restriction into some numerical examples in Section 4.
 As noted above, when $\theta=1$, the best $n$ is $n=1$.
-In that case RQMC is the same as MC, using one uniformly distributed point in $[0,1]$.
+In that case RQMC is the same as MC, using one uniformly distributed point in $0,1]^{d}$.
 This result holds for any sampler that gets the ordinary MC rate, even if it has a favorable constant.
@@
 We begin with finite $N$ and for some integrands with known RQMC variance, we find the $n$ that minimizes the interval widths from Bennett’s inequality.
-They are then the exact best $n$ for Bennett’s inequality which the HBCIs and EBCIs of WauRam24a match asymptotically.
+They are then the exact best $n$ for Bennett’s inequality which the HBCIs and EBCIs of Waubdy-Smith and Ramdas (2024) match asymptotically.
 After that we present finite sample results on the EBCIs and HBCIs for a variety of integrands differing in smoothness and dimension.
@@
 For our HBCI computations, we used $\theta=1/2$.
-The software we used comes from hedged_algo which has code to reproduce the figures in WauRam24a .
-It calls the HBCI and predictable plug-in EBCI algorithms from bet_algo .
-The scrambled Sobol’ points we used were from QMCPy QMCPy2020a .
+The simulations presented here were run at Betting-Paper-Simulations-for-QMC.
+The software we used comes from Waudby-Smith (2025) which has code to reproduce the figures in Waubdy-Smith and Ramdas (2024).
+It calls the HBCI and predictable plug-in EBCI algorithms from Howard et al.
+(2021).
+The scrambled Sobol’ points we used were from QMCPy Choi et al.
+(2025).
 4.1 Semi-empirical evaluations
@@
 We can compute these confidence interval widths for very large values of $N$ that would be impractical to simulate.
-The digital nets of Sobol’ Sob67 under the nested uniform scramble of Owe95 have a stratification property.
-The $n$ values of $x_{ij}$ for $i=1,\dots,n$ and any component $j=1,\dots,d$ have one value uniformly distributed in each interval $[(\ell-1)/n,\ell/n)$ for $\ell=1,\dots,n$ and those $n$ values are mutually independent.
+The digital nets of Sobol’ Sobol’ (1967) under the nested uniform scramble of Owen (1995) have a stratification property.
+The $n$ values of $x_{ij}$ for $i=1,\dots,n$ and any component $j=1,\dots,d$ have one value uniformly distributed in each interval $(\ell-1)/n,\ell/n)$ for $\ell=1,\dots,n$ and those $n$ values are mutually independent.
 This allows us to work out the finite sample RQMC variances for some illustrative examples.
-The random linear scramble of Mat98 combined with a digital shift attains the same variance as the nested uniform scramble, but it has some dependencies among the stratified sample values.
+The random linear scramble of Matoušek (1998) combined with a digital shift attains the same variance as the nested uniform scramble, but it has some dependencies among the stratified sample values.
 In both cases, the $d$ vectors $(x_{1j},\dots,x_{nj})\in\mathbb{R}^{n}$ are mutually independent.
-We only consider values of $n$ that are powers of $2$ which is a best practice when using Sobol’ sequences Owe22a .
+We only consider values of $n$ that are powers of $2$ which is a best practice when using Sobol’ sequences Owen (2022).
 The smooth error rate $\tilde{O}(n^{-3/2})$ for the RMSE is attainable along powers of $2$ but not for general $n$.
-An elegant argument due to Sobol’ sobo:1998 explains why.
+An elegant argument due to Sobol’ Sobol (1998) explains why.
 He notes that
@@
 That integrand is $f(x)=1\{x<1/3\}$.
-It is constant within $n-1$ of the intervals $[(\ell-1)/n,\ell/n)$.
+It is constant within n1n-1 of the intervals $(\ell-1)/n,\ell/n)$.
 All of this function’s variation happens within just one of those intervals.
-By inspecting the base $2$ expansion $1/3=0.010101\cdots$ we can see that the remaining value of $f(x_{\ell})$ is $1$ or 00 and it equals $1$ with a probability that is either $1/3$ if $\log_{2}(n)$ is an odd integer or is $2/3$ if $\log_{2}(n)$ is an even integer.
+By inspecting the base $2$ expansion $1/3=0.010101\cdots$ we can see that the remaining value of $f(x_{\ell})$ is $1$ or 0 and it equals $1$ with a probability that is either $1/3$ if $\log_{2}(n)$ is an odd integer or is $2/3$ if $\log_{2}(n)$ is an even integer.
 It follows that $\operatorname{var}(\hat{\mu}_{n})=2/(9n^{2})$.
@@
 The value $N^{2/3}W$ very quickly becomes constant.
-For non-integer $\theta$ (not shown here) the scaled widths fluctuation within a finite range instead of approaching a constant.
+For non-integer $\theta$ (not shown here) the scaled widths fluctuate within a finite range instead of approaching a constant.
 The result in Table 1 for $N=1024$ is to choose $n=8$ which violates the guidance that the best $n$ is at most $7$ when $N=1024$ and $\alpha=0.05$.
 The reason is that this table restricts to powers of $2$, and $n=8$ is better than $n=4$.
-$\log_{2}(N)$ $N$ $n$ $W$ $N^{2/3}W$
-00 1 1 5.02e$+$00 5.020114
-04 16 2 7.60e--01 4.826379
-07 128 4 1.90e--01 4.826379
-10 1,024 8 4.75e--02 4.826379
-13 8,192 16 1.19e--02 4.826379
-16 65,536 32 2.97e--03 4.826379
-19 524,288 64 7.42e--04 4.826379
-22 4,194,304 128 1.86e--04 4.826379
-25 33,554,432 256 4.64e--05 4.826379
-28 268,435,456 512 1.16e--05 4.826379
-Table 1: As $N$ varies for $f(x)=1\{x<1/3\}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{2/3}W$.
+$\log_{2}(N)$ NN $n$ $W$ $N^{2/3}W$
+00 1 1 5.02e++00 5.020114
+04 16 2 7.60e-01 4.826379
+07 128 4 1.90e-01 4.826379
+10 1,024 8 4.75e-02 4.826379
+13 8,192 16 1.19e-02 4.826379
+16 65,536 32 2.97e-03 4.826379
+19 524,288 64 7.42e-04 4.826379
+22 4,194,304 128 1.86e-04 4.826379
+25 33,554,432 256 4.64e-05 4.826379
+28 268,435,456 512 1.16e-05 4.826379
+Table 1: As NN varies for $f(x)=1\{x<1/3\}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{2/3}W$.
 Each row corresponds to a value of $N$ for which $n$ has increased.
 Our second example is $f(x)=xe^{x-1}$.
-This function is bounded between 00 and $1$.
+This function is bounded between 0 and $1$.
 We chose it because it is smooth enough to have $\theta=3$ while also not a polynomial or an antisymmetric function which would make the problem artificially easy for some methods.
@@
 Table 2 shows how the optimal sample sizes $n$ grows with $N$ for our smooth example $f(x)=xe^{x-1}$.
-An RQMC sample size of $n=8$ becomes optimal for $N=4{,}096$ and remains optimal up to $N=32{,}768$.
+An RQMC sample size of $n=8$ becomes optimal for N=4,096N=4{,}096 and remains optimal up to N=32,768N=32{,}768.
 The values $N^{3/4}W$ settle down rapidly as $N$ increases.
 Consistent with the results of Section 3, the values $n$ grow at a slower rate with $N$ than in the BVHK case while the widths $W$ decrease at a faster rate.
-$\log_{2}(N)$ $N$ $n$ $W$ $N^{3/4}W$
-0 1 1 4.00e$+$00 4.003728
-4 16 2 5.17e--01 4.138086
-8 256 4 6.52e--02 4.175717
-12 4,096 8 8.17e--03 4.185407
-16 65,536 16 1.02e--03 4.187848
-20 1,048,576 32 1.28e--04 4.188459
-24 16,777,216 64 1.60e--05 4.188612
-28 268,435,456 128 2.00e--06 4.188651
-Table 2: As $N$ varies for $f(x)=xe^{x-1}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{3/4}W$.
+$\log_{2}(N)$ NN $n$ $W$ $N^{3/4}W$
+0 1 1 4.00e++00 4.003728
+4 16 2 5.17e-01 4.138086
+8 256 4 6.52e-02 4.175717
+12 4,096 8 8.17e-03 4.185407
+16 65,536 16 1.02e-03 4.187848
+20 1,048,576 32 1.28e-04 4.188459
+24 16,777,216 64 1.60e-05 4.188612
+28 268,435,456 128 2.00e-06 4.188651
+Table 2: As NN varies for $f(x)=xe^{x-1}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{3/4}W$.
 Each row corresponds to a value of $N$ for which $n$ has increased.
@@
 $g_{\mathrm{finance}}(v)$ $=\min\Bigl(1,\frac{\sqrt{\max(v+2,0)}}{2}\Bigr).$
-All of the above choices give ridge functions bounded between 00 and $1$.
+All of the above choices give ridge functions bounded between 0 and $1$.
 They span a very wide range of QMC regularity conditions.
 Using $g_{\mathrm{jump}}$ makes $f$ discontinuous and it has infinite variation in the sense of Hardy and Krause for $d\geqslant 2$.
-Using $g_{\mathrm{kink}}$ makes $f$ have infinite variation in the sense of Hardy and Krause for $d\geqslant 3$ and yet it has been seen to be a more ‘QMC friendly’ integrand than $g_{\mathrm{jump}}$ in high dimensions because it has an asymptotically bounded mean dimension mdridge , meaning that it is dominated by low dimensional interactions.
+Using $g_{\mathrm{kink}}$ makes $f$ have infinite variation in the sense of Hardy and Krause for $d\geqslant 3$ and yet it has been seen to be a more ‘QMC friendly’ integrand than $g_{\mathrm{jump}}$ in high dimensions because it has an asymptotically bounded mean dimension Hoyt and Owen (2020), meaning that it is dominated by low dimensional interactions.
 It is continuous, has one weak derivative but is not differentiable everywhere.
 Using $g_{\mathrm{smooth}}$ makes $f$ infinitely differentiable which is generally a very QMC friendly property.
-Using $g_{\mathrm{finance}}$ gives $f$ a property similar to some integrands in financial option valuation, namely an unbounded derivative where $v=-2$.
+Using $g_{\mathrm{finance}}$ gives $f$ a property similar to some integrands in financial option valuation, namely an unbounded derivative where v=2v=-2.
 For each of these four ridge functions, we considered budgets $N=2^{K}$ for $K\in\{8,10,12,14,16\}$ with RQMC sample sizes $n=2^{k}$ for $k\in\{0,1,2,3,4,5,6\}$ in dimensions $d\in\{1,2,4,16\}$.
@@
 Figure 1: Mean confidence interval width for the HBCI as a function of the RQMC sample size $n$.
-There is one panel for each of the $4$ ridge functions.
+There is one panel for each of the 44 ridge functions.
 Within a panel the curves are for $N=2^{2r}$ for $r=4,5,6,7,8$ from top to bottom.
-The smallest mean width for each $N$ is marked with a larger circle.
-The means are taken over all $20$ replicates and all $4$ dimensions used.
+The smallest mean width for each NN is marked with a larger circle.
+The means are taken over all 2020 replicates and all $4$ dimensions used.
 We see in Figure 1 some things that align very well with our analysis and some that were not anticipated.
@@
 We know that the HBCI have lengths that asymptotically match Bennett’s formula, but we also want to study how those lengths approach Bennett’s formula in some examples.
-Using the known variances, we can compute the half width from Bennett’s inequality  (11) for our $\sigma^{2}_{n}$ and $R=N/n$ double it and then divide the average HBCI width of 20 independent trials by this value.
+Using the known variances, we can compute the half width from Bennett’s inequality (11) for our $\sigma^{2}_{n}$ and $R=N/n$, double it, and then divide the average HBCI width of 20 independent trials by this value.
 The results are shown in Figure 2.
@@
 $1\{x<1/3\}$ $xe^{x-1}$
-$N$ $n_{*}$ $n_{\mathrm{opt}}$ $n_{*}$ $n_{\mathrm{opt}}$
+NN $n_{*}$ $_{\mathrm{opt}} nn_{*}$ $n_{\mathrm{opt}}$
 256 4 4 4 2
@@
 65,536 32 16 16 8
-Table 3: For the given budgets $N$, we show $n_{*}$ as the minimizer of the Bennett width (13) over $n=2^{k}$ for integer $k$ and $n_{\operatorname{opt}}$ is the value of $n$ that minimized the average width of the HBCI for the given $N$.
+Table 3: For the given budgets NN, we show $n_{*}$ as the minimizer of the Bennett width (13) over $n=2^{k}$ for integer $k$ and $n_{\operatorname{opt}}$ is the value of $n$ that minimized the average width of the HBCI for the given NN.
 These are shown for functions $f(x)=1\{x<1/3\}$ and $f(x)=xe^{x-1}$.
@@
 In favorable cases the RQMC variance is $\tilde{O}(n^{-3})$ and then using a fixed number $R$ of replicates of point sets with $n\to\infty$ we can estimate the integral $\mu$ with a standard deviation of $\tilde{O}(N^{-3/2})$ for $N=nR$ along with an unbiased estimate of the variance of that estimate.
-It remains very difficult to get a confidence interval of width $\tilde{O}(N^{-3/2})$ err4qmc .
-Using Student’s $t$ with a small $R$ gave good results in an extensive simulation by LEcEtal24a but that success is not yet theoretically understood.
+It remains very difficult to get a confidence interval of width $\tilde{O}(N^{-3/2})$ Owen (2026).
+Using Student’s $t$ with a small $R$ gave good results in an extensive simulation by L’Ecuyer et al.
+(2024) but that success is not yet theoretically understood.
 For an integrand subject to known bounds, we can get a non-asymptotic confidence interval by using $R$ replicates of $n$ RQMC points using a predictable plug-in empirical Berstein confidence interval.
@@
 In a task like that it makes sense to only use $n=2^{k}$ for integers $k$.
-That is partly because powers of two are good for scrambled Sobol’ points, but also because in RQMC it is generally advisable to study sample sizes $n$ that grow geometrically not arithmetically sobo:1998 .
+That is partly because powers of two are good for scrambled Sobol’ points, but also because in RQMC it is generally advisable to study sample sizes $n$ that grow geometrically not arithmetically Sobol (1998).
 The idea is that if $n$ is not large enough to get a good answer, then $n+1$ is unlikely to be much of an improvement.
@@
 For scrambled Sobol’ points it is strongly advisable to take $n=2^{k}$ because those are the best sample sizes and they can even give a better convergence rate.
-The Halton sequence Hal60 does not have strongly superior sample sizes and so $n$ need not be a power of two for it.
-The two scrambles we considered for Sobol’ points can also be applied to the Halton sequence getting $\operatorname{var}(\hat{\mu})=o(1/n)$ without requiring special sample sizes $n$ haltongain .
+The Halton sequence Halton (1960) does not have strongly superior sample sizes and so $n$ need not be a power of two for it.
+The two scrambles we considered for Sobol’ points can also be applied to the Halton sequence getting $\operatorname{var}(\hat{\mu})=o(1/n)$ without requiring special sample sizes $n$ Owen and Pan (2024).
 There are however no published non-trivial settings where scrambled Halton points can obtain $\theta>2$.
@@
 We thank Aaditya Ramdas and Ian Waudby-Smith for helpful discussions.
+We thank an anonymous reviewer for helpful comments.
 ABO was supported by the U.S.